### v3

# Right Place, Right Time! Dynamizing Topological Graphs for Embodied Navigation

Abstract
Embodied Navigation tasks often involve constructing topological graphs of a scene during exploration to facilitate high-level planning and decision-making for execution in continuous environments. Prior literature makes the assumption of static graphs with stationary targets, which does not hold in many real-world environments with moving objects. To address this, we present a novel formulation generalizing navigation to dynamic environments by introducing structured object transitions to dynamize static topological graphs called Object Transition Graphs (OTGs). OTGs simulate portable targets following structured routes inspired by human habits. We apply this technique to Matterport3D (MP3D), a popular simulator for evaluating embodied tasks. On these dynamized OTGs, we establish a navigation benchmark by evaluating Oracle-based, Reinforcement Learning, and Large Language Model (LLM)-based approaches on a multi-object finding task. Further, we quantify agent adaptability, and make key inferences such as agents employing learned decision-making strategies generalize better than those relying on privileged oracle knowledge. To the best of our knowledge, ours is the first work to introduce structured temporal dynamism on topological graphs for studying generalist embodied navigation policies. The code and dataset for our OTGs will be made publicly available to foster research on embodied navigation in dynamic scenes.
Figure 1: Object Transition Graph (OTGs): We introduce dynamism to topological graphs via portable targets following structured object transition scenarios. In this representative figure, the watch (in red) moves from node N1 at T=10:00 AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM. We study the performance of various navigation agents in different object transition scenarios, and establish their generalizability.
1 Introduction
Embodied navigation tasks often rely on topological graphs (TGs) for decision-making. Visual-Language Navigation (VLN) tasks such as REVERIE [36] and R2R [2, 11] utilize TGs extracted from MP3D [8] to follow human guidance. Similarly, methods for ObjectNav, PointNav, and ImageNav often employ TGs  [20, 53, 14, 55, 10, 22, 38], where nodes represent locations or landmarks, and edges define traversable paths. These graphs provide a structured way to plan and execute navigation policies.
A key limitation of these existing embodied navigation paradigms is their reliance on static graphs, where objects within nodes remain stationary. This static assumption is unrealistic, as real-world environments are inherently dynamic, where users frequently move small, portable objects from place to place, such as shifting their phone or wallet between rooms throughout the day. Studies on robots in collaborative environments  [34, 3, 16] suggest that human object placements follow structured patterns shaped by routines and habits, introducing a degree of entropy in object locations over time. Despite this, existing navigation benchmarks fail to model such temporal variability, limiting their applicability for real-world deployment.
In the ObjectNav task [18, 4] for instance, common state-of-the-art navigation approaches involve Reinforcement Learning (RL) [54, 52, 4, 47] and LLM-based zero-shot methods [18, 45, 6] to find stationary objects. RL policies for multi-object settings [50, 9] have also been introduced, but these are also based on static environments. For real-world deployment, these approaches must generalize, or maintain performance in dynamic scenes with shifting targets. Generalizability in prior embodied work has been described in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to static environments that were unseen or used synthesized guidance language [57, 26, 23, 17]. However, there has been little to no work that defines generalizability in the context of adapting to dynamic settings with moving objects. The evaluation of these prior approaches on dynamic settings with non-stationary objects remains to be seen.
To address this gap, in this work, we introduce Object-Transition Graphs (OTGs), a dynamic variant of topological graphs that incorporates structured object transitions over time. Figure 1 illustrates this concept. Unlike traditional TGs, where navigation is performed on a static graph, OTGs introduce a temporal component, where objects shift positions while the underlying graph structure remains constant. This forces an agent to synchronize its trajectory with temporal object movements; its adaptability to change allows us to measure its generalist navigation performance. In dynamizing static topological graphs, our work seeks to redefine how we evaluate decision-making on embodied agents, by emphasizing not just where to navigate but also when.
In this work, we define the generalizability of a navigation approach by comparing the performance between static TGs and their OTG counterparts. This allows us to quantify an agent’s robustness to dynamic changes, providing us with key insights on the effectiveness of graph-based navigation schemes in real-world settings.
Main Results. We generalize embodied navigation to dynamic environments by 1) introducing structured object transitions to make static topological graphs dynamic, and 2) presenting a benchmark for evaluating the adaptability of state-of-the-art navigation policies in such dynamic settings. Our contributions are summarized as follows:
  • Dynamizing Topological Graphs: We propose a novel framework for introducing structured object transitions on static topological graphs, transforming them into Object Transition Graphs or OTGs. Inspired by human habits, our approach defines three distinct object transition scenarios to govern the movement of small, portable objects in a scene, providing varying levels of entropy. Our method applies to any static topological graph, including those obtained from Matterport3D (MP3D) [8] and HM3D [37]. Our dynamized MP3D environments and code will be released as an open-source benchmark to foster research in this area.
  • OTG Benchmark: We introduce a benchmark for evaluating navigation performance on OTGs, comparing heuristic, RL, and LLM-based agents across structured object transition scenarios on a multi-object finding task. Measuring performance of agents on OTGs provides insights into the adaptability of graph-based approaches in dynamic environments with non-stationary objects.
  • Generalization in Dynamic Environments: We make several key inferences on the generalizability of various navigation approaches by comparing their performance on OTGs against static TGs. In particular, we observe that strategies relying on privileged knowledge can lead to suboptimal decision-making in dynamic environments, while approaches that improve agents with experience and collected observations better generalize to OTGs.
2 Related Works
Figure 2: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table 1) and transit scenarios (Table 2). Here, a mug is placed on a kitchen node $N_{3}$ at timestep $T=53$, but moves to a bedroom node $N_{5}$ at timestep $T=55$ via node $N_{4}$. If an agent reaches the kitchen after $T=53$ (or the bedroom before $T=55$), it would fail to see the mug. Multiple objects could also be at the same node (hat and mug in $N4$).
Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [41, 31, 5] and socially aware navigation [7, 29], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time. In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments, with limited research on handling dynamic targets  [11, 44]. With OTGs, we introduce a form of structured temporality in dynamic environments where agents must reason over spatio-temporal changes, extending high-level planning to dynamic scenes with shifting targets. While recent simulators like Habitat 3.0 [35] include object rearrangement tasks, these primarily focus on agents modifying the environment rather than adapting to dynamic object placements over time.
Planning in dynamic environments has been studied in the past [48, 59, 32], with recent approaches even utilizing LLMs in conjunction with multi-arm bandits [15]. These schemes usually propose memory augmentations with hierarchical planning procedures and frame the problem from an obstacle avoidance standpoint [48, 25, 58]. In contrast, our work considers the case where the objects themselves are non-stationary.
Rudra et al. [40] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints. In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to an object-finding navigation task in a static environment. In contrast, we tackle a truly dynamic case, where objects are moving even during the episode. This definition adds a layer of complexity as the embodied agent must now navigate towards a constantly shifting target object, for which it needs to identify specific routines and object movement patterns in the environment.
Kurenkov et al. [21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations. They also performed target object-finding experiments, but the environment was static as the agent moved. In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs. Furthermore, Wang et al. [49] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments. In contrast, our work focuses on converting topological graphs into object transition graphs, and comments on the generalizability of embodied navigation approaches in dynamic environments.
3 Object Transition Graph (OTG):
We define Object Transition Graphs (OTGs) is an evolutionary undirected graph where the underlying graph structure remains constant, but the objects present in its nodes evolve over time. Formally, let $G=(V,E)$ be a topological graph structure with nodes $V$ representing spatial locations (rooms and objects), and edges $E$ representing traversable paths between the nodes. Let $\mathcal{O}$ represent the total set of available portable objects. Then, at each timestep $t$, $v\in V$ is associated with a set of portable target objects $O_{v}(t)\subseteq\mathcal{O}$. While the graph topology remains static, the object assignments $O_{v}(t)$ evolve via a time-dependent transitioning style $\Lambda$, to give us an object transition function $O$ as,
$O:V\times\Lambda\rightarrow 2^{\mathcal{O}}$ (1)
where $2^{\mathcal{O}}$ is a power set representing all possible portable object combinations. At each timestep $t$, $O(v,t)$ gives us the portable targets present in each node $v$ based on the transition style $\Lambda$.
Room Portable Objects
Bedroom Charger, Water Bottle, Smartwatch, Laptop, Notebook, Toothbrush, Mug, USB Flash Drive, Phone, Headphones, Hat
Garage Screwdriver, Flashlight, Mug, Phone, Headphones, Hat
Dining Salt and Pepper Shakers, Portable Speaker, Charger, Water Bottle, Mug, Bowl, Phone, Headphones, Hat
Office Charger, Laptop, Hat, Notebook, USB Flash Drive, Mug, Phone, Headphones
Bathroom Toothbrush, Phone, First-Aid Kit
Kitchen Salt and Pepper Shakers, Hat, Mug, Bowl, Phone, Headphones, First-Aid Kit
Lounge Playing Cards, Mug, Portable Speaker, Charger, Water Bottle, Laptop, Phone, USB Flash Drive, Dice, Headphones, Hat
Gym Dumbbells, Jumprope, Smartwatch, Phone, Headphones, Hat
Outdoor Jumprope, Smartwatch, Portable Speaker, Phone, Water Bottle, Headphones, Hat
Recreation Playing Cards, Dice, Water Bottle, Headphones, Hat
Table 1: Rooms and Portable Objects: We map $21$ portable objects to a set of household rooms. This mapping is used to set destinations for object transit. During each episode, objects are placed in various rooms for a range of timesteps. Commonly moved objects such as phone, headphones are associated with $9$ different rooms, while less commonly shifted ones such as dumbbells appear only in the Gym.
Algorithm 1 Modify $G$ With Object Transitions $\Lambda$
Input: Topological Graph $G=(V,E)$, Portable Object Set $\mathcal{O}$, Episode Length $T$, Initial Agent Node $r$, Object Density $d$, Object Transit Scenario $\Lambda$, Local Graph Size $q$.
Output: Evolving Graph Trajectory $\zeta$
Initialize: Add $\mathcal{O}$ portable objects to $G_{0}$. Calculate the object density $d$ of the local graph $g$ around the initial node $r$. $g$ is all nodes within $q$ steps of $r$.
Sample $N\in\mathcal{O}$ duplicate new objects and add them to random nodes in $g$ to match required density $d$. $\mathcal{O}^{\prime}=\mathcal{O}+N$ is the total set of objects at $G_{0}$
1:  for $t\in[1,\dots,T]$ do
2:     $G_{t}=G_{t-1}$
3:     for each object $o\in\mathcal{O}^{\prime}$ do
4:        Using Equation 1: $O(v,t)\leftarrow O(v,\Lambda)$
5:        Sample $v_{k}\in V$ from $O(v,t)$Following $\Lambda$}
6:        if $v_{k}\neq v_{o}$ then
[…]
8:              Use fixed trajectory $\delta(t)$
9:           else
10:              Compute random path $\delta(t)$ from $v_{o}$ to $v_{k}$
11:           end if
12:           Move $o$ along $\delta$ to $v_{k}$ over $t$ timesteps
13:        end if
14:     end for
15:     Add $G_{t}$ to $\zeta$
16:  end for
17:  return $\zeta$
3.1 Defining Transition Scenarios $\Lambda$:
Scenario ($\mathbf{\Lambda}$) Fixed Rooms Fixed Paths Entropy
Static N/A None
Random High
Semi-Routine Medium
Fully-Routine Low
Table 2: Object Transit Scenarios $\mathbf{\Lambda}$: Portable targets transition in the graph under structured scenarios inspired by human habits. In the random case, the portable objects can move to any room at any time during each episode. In the routine cases, the rooms that the target objects can travel to are fixed (except during transit). The static scenario is a baseline with stationary items i.e., zero entropy.
Human habit formation via object placement can be attributed to cognitive offloading, where individuals rely on the environment to reduce memory demands [39, 46]. This reliance on the environment influences the spatial distribution of objects. Frequently used objects (like keys or phones) might move more flexibly across rooms (high entropy), while function-specific objects (like dumbbells or toothbrush) might be more constrained (low entropy). We use this structured transition behavior to define an object’s transition entropy, which we use to formulate $\Lambda$. We consider three object transit scenarios at decreasing levels of transit entropy:- 1) Random, Semi-Routine, and Fully-Routine. These are outlined in Table 2.
(a) Random Scenario
(b) Semi-Routine Scenario
(c) Fully-Routine Scenario
Figure 3: Object Frequency Visualization: A visual representation of portable target frequencies over several timesteps in various transit scenarios $\Lambda$. We simulate portable objects transiting on a topological graph from MP3D. In the random case, object frequencies are more scattered (high entropy); in semi-routine, target rooms are fixed, but paths are flexible (medium entropy); and in the routine case, both rooms and paths are fixed (lowest entropy). Note the distribution of object frequencies visually correlates to their entropy.
Algorithm 1 describes our approach for generating an evolving graph trajectory $\zeta$ to dynamize any topological graph $G$. Figure 2 further illustrates this setup. The initial graph contains portable objects placed according to Table 1. We then generate $\zeta$ by moving objects in the graph at various timesteps for each transit scenario $\Lambda$. $\zeta$ is precomputed for each transit scenario and stored for later analysis with navigation algorithms.
When generating the evolving graph, a pseudo-random seed $s$ helps differentiate the movement of portable objects per episode. Letting $e$ be the episode index,
$s=\begin{cases}1\>\text{(or fixed)}&\text{if movement is fully routine},\\ e&\text{if movement is random or semi-routine}.\end{cases}$ (2)
Note that for the fully routine case, $\zeta$ remains constant for all episodes of the experiment, meaning that the object transition routes remain fixed. In the other two cases, $\zeta$ varies per episode according to the room and interval variability described. Figure 3 presents a visual representation of the object frequencies in all three scenarios.
For our experiments, we dynamize topological graphs obtained from the Matterport3D (MP3D) [8] dataset using Algorithm 1 to obtain OTGs. Each node contains panoramic images representing the scene, and edges represent the distance between them. Additionally, we have the room type for each node and use them as a template to define the set of portable objects listed in Table 1.
Note on Visual Spatial Placement. Beyond object transitions, we also investigate spatial transformations in placing the objects on the scene. Portable objects are small and can be pasted on the scene in various ways. We first use FastSAM [56] to obtain segmentation masks for both the scene and the portable target object. On the scene segmentations, we filter out a set of $K$ largest segments below the center of the image, to avoid placing floating targets. We then place the target objects on a randomly chosen segment $k\in K$ this set. This simple placement technique for small objects allows us to modify images of pre-existing topological graphs and makes the visual grounding task challenging enough for comparison against various backgrounds. While 3D object placement would be more accurate, this is more cumbersome to setup, and is tangential to the focus of this paper, which is to introduce OTGs to study navigation generalizability. More details on our visual grounding experiments with VLMs can be found in the Appendix. We will be releasing this dataset of small objects placed on MP3D to foster research in embodied visual grounding.
4 Task: Multi-Object Finding
To evaluate how well different navigation methods perform on dynamic OTGs, we aim to assess their generalizability. Given a topological graph, the generalizability is the difference in performance achieved by a navigation scheme between when then graph is static and when the graph is dynamic. For example in Section 6, we measure the generalizability of PPO-based agents by training an agent in a static environment and training another in a dynamic environment and compare the difference in performance.
To enable this study, we focus on the task of finding multiple portable target objects over a fixed timespan. This setting adds realism to other multi-object finding tasks such as Multi-ON [50] and GOAT [9] by introducing non-stationary targets. For dynamic settings such as ours, we argue that a multi-object finding task is preferable over a standard single-object finding setup for two reasons:-
  • Realistic Setting: In real-world homes, users frequently move objects, leading to dynamically changing object locations [3, 34]. Assistive agents must track and infer object movements over time to provide meaningful assistance. This requires monitoring multiple portable objects (e.g., phones, watches) rather than treating each search as an isolated event.
  • Lifelong Navigation: A multi-object setup enables an agent to finetune its search policy in real-time, by using knowledge gathered from finding previous objects to better find the next. An agent deployed at homes will likely have to face similar conditions, with a user asking it to find multiple targets without ‘resetting’ its position, much akin to lifelong navigation [53].
First, we mathematically formalize what navigating on an OTG entails. Let $\Lambda$ describe the transitory motions of objects. Then, let $O$ be the set of portable objects found over $T$ timesteps for an agent starting at node $r$ with an object distribution density $d$. Then $S=[r,d,T,\Lambda]$ represents a set of our experimental variables for this task. We then define our navigation objective as “Finding the maximum number of portable objects $O$ while traversing a dynamic environment represented by $S$.”.
For downstream applications like embodied agents personalizing to households, we believe finding multiple targets to be a good task to study dynamic adaptability, due to dense reward signals [9]. The setting is also realistic as households have multiple people moving objects around at the same time. We study the performance of different navigation policies with varying simulator conditions $S$.
In each episode, the agent is tasked with finding as many portable target objects as possible by making decisions on a dynamic OTG. Let $\tau$ be a finite trajectory representing a sequential list of visited nodes on the OTG, $G$, and let $\tau(t)$ be the $t$-th node in $\tau$. We can now define $O(\tau,t)$ as the set of portable objects found at the $t$-th node of $\tau$ at time $t$. Given that the objects move throughout the scene, the $t$-th node may have a different set of portable objects present at a different timestep. Given a trajectory $\tau$ of length $T$, we can formulate the total set of portable objects found along $\tau$ as $O_{\tau}=\cup_{t=1}^{T}O(\tau,t)$.
A navigation agent parameterized by policy $\pi$ explores the dynamic OTG by generating a trajectory $\tau_{\pi}$, where each timestep $t$ corresponds to a selected node: $\tau_{\pi}(t)\in V$. Specifically, $\tau_{\pi}(t)$ is the node chosen at time $t$ by policy $\pi$. We now can write the policy optimization problem as finding an optimal policy $\pi^{*}$ defined as $\pi^{*}=\operatorname*{arg\,max}_{\pi}|O_{\tau_{\pi}}|$.
Consequently, $O_{\tau_{\pi^{*}}}$ is the maximum number of portable targets that can be found by taking the optimal trajectory. In the following section, we describe various navigation schemes that we employ to perform the multi-object finding task on OTGs.
5 Navigation Approaches
To infer the generalizability of prior navigation models, we emulate a static topological graph environment by keeping portable objects stationary in one location throughout the experiment. We then evaluate both the static topological graph environment and the dynamic OTG environment with $5$ relevant navigation agents.
Random Agent. For a completely random navigation baseline, at each timestep $t$, the agent chooses one of its neighboring nodes to go to for the next iteration.
Oracle Agent. We utilize a greedy heuristic on an agent as an oracle baseline: At each timestep $t\in[1,T]$, we perform a Breadth-First Search (BFS) to find the closest node with an unseen portable object. Using the shortest path between the unseen object and its current position, the agent then moves its neighboring node on the given shortest path. This approach relies on the agent having access to an “oracle” that tells it about the closest node containing a unique portable target at a given timestep. While oracle knowledge is useful since the position of the objects in the environment changes with time, note that there is a possibility that it will not find an object after navigating to an oracle-guided node.
Reinforcement Learning Agent. We use a reinforcement learning-based method, specifically Proximal Policy Optimization (PPO) [43] to define an agent. PPO has been used in several previous works in embodied navigation tasks, such as ObjectNav [54, 60, 51] with SOTA results.
Observation Space: Given the limited prior literature dealing with PPO on a dynamic topological graph, we consider constructing our agent with the following observation space as input to the navigation agent. To define the observation space, let $|G|$ be the number of nodes on $G$, and each node is given a unique ID from $\{0,\dots,|G|-1\}$:
  • Current timestep, $t\in T$
  • ID the current node, $n_{t}$ from a list of all nodes in the environment graph
  • A list with each index corresponding to a node ID. The element at a given index is the number of objects at the corresponding node at timestep $t$
Action Space: The PPO agent then outputs the index of the node it wants to travel to for the next timestep $t+1$. We use a Maskable PPO to only allow it to choose neighboring nodes. The mask in this case is a list with each index corresponding to a node ID. The element at a given index is either $1$ if it is a neighbor of the current node and 00 otherwise.
Reward Objective: Let $O_{t-1}$ be the set of distinct portable objects the agent as seen up to time $t-1$. The reward is the number of new objects it has found at the current node $n_{t}$ that are not in $O_{t-1}$. Formally, the reward at timestep $t$ is $|O(n_{t})\setminus O_{t-1}|$.
Note. Our version of PPO uses a map of the environment and the number of objects at each node at each timestep.
LLM-based Agents. For LLM-based navigation, we consider two variations of LGX [18], an LLM-based method for zero-shot embodied exploration.
1. LLM-Vanilla: Given a target object, the LLM takes a list of objects around the agent and asks an LLM (GPT-4o) to predict an object to navigate toward. The agent then navigates towards the predicted object and continues doing so till the provided target is found.
The LLM’s prompt contains a set of portable targets remaining to be found in the environment. At each timestep, we use YOLO v8 [19] to obtain a list of objects in the scene, and then prompt GPT-4o with the object list asking for an object prediction from the list that might lead to an unseen portable object. We then use the MP3D topological graph to hop to the next node.
2. LLM with Memory: The vanilla version of the LLM agent works in a zero-shot manner resembling a Markov Chain, meaning that it makes navigation decisions at each node solely based on the observations it has seen at the current timestep. Since objects keep transiting on our OTG, keeping track of previously seen objects should help the agent make more prudent navigational decisions.
We accommodate this by incorporating memory by passing a set of historical objects seen, predictions made, and timesteps to the prompt across episodes. The system prompt contains this historical data, along with a set of portable objects remaining to be found in the environment. We maintain a horizon to avoid GPT token overflow, where we remove the earliest appended observation when an overflow of tokens occurs.
6 Experiments and Results
We treat navigating in dynamic scenes and visual target grounding as separate problems in our experiments. This decoupling allows us to individually analyze navigation and grounding performance. The results of our visual grounding experiments are explained in detail in the supplementary.
6.1 Hyperparameters
We determine 4 hyperparameters that influence performance as described in the definition section, i.e., $S=[r,d,T,\Lambda]$. Each agent is subjected to a total of $[Stationary(1)+Transit(3)]\times R\times E\times T$ experiments, where $R$ is a set of rooms that varies with each scan from which a random node $r\in R$ is picked as a starting point for each trial. The maximum number of trials is thus equal to the number of rooms. $R$ ranges from $[11,30]$ in our chosen subset of $10$ MP3D scans. $E$ is the number of episodes that we run for each starting node $r$, and $T$ is the number of timesteps per episode. We set $R=10$, $E=20$ and $T=30$ in our experiments. Although different objects can have different levels of transition entropy, we set all objects in an experimental run to have the same transit scenario $\Lambda$. Setting all objects to have the same transit allows us to better analyze the generalizability of a navigation scheme in different levels of entropy in the OTG.
Ground Truth Paths ($\Pi^{*}$). Given a starting node $r\in R$, for each episode $e\in E$, there exists a set of ground truth paths $\Pi^{*}$ over $T$ that the agent could take for collecting the most number of portable targets. We compute these paths by simulating all possible trajectories from a given starting point $r$, and store $\Pi^{*}$ to calculate performance and efficiency later.
Object Density $d$. To make the MP3D environment less sparse with objects, we define a local graph of all nodes within $3$ steps away from the agent’s initial node. We then populate the local graph by adding duplicate portable objects from $O$ to the local graph. We populate until the object density $d$ of this local graph is $1$. As in, on average, each node in the local graph has $1$ object. In Algorithm 1, $d$ and the local graph size $q$ are hyperparameters.
6.2 Evaluation Metrics
Success Rate (SR): Given an episode, we measure the object-finding performance of an agent by dividing the number of objects found in that episode by the maximum possible number of objects that could have been found with a ground-truth path $\pi^{*}\in\Pi^{*}$: $SR=|O_{\pi}|/|O_{\pi^{*}}|$. This metric is similar to the Progress metric in Wani et al. [50]. In this work, the ground-truth trajectory may not necessarily find all the objects in the given timeframe.
Trajectory Alignment (TA): To measure path efficiency, we define Trajectory Alignment as the average overlap between ground-truth trajectory $\pi^{*}$ and trajectory taken by the agent across episodes. We take the maximum across $\Pi^{*}$: $TA=\min_{\pi^{*}\in\Pi^{*}}\frac{1}{T}\sum_{t\in T}\mathbf{1}_{\pi(t)=\pi^{*}(t)}$, where $\mathbf{1}$ is an indicator function.
Relative Change in Success (RCS): We use this value taken from [17] to measure the generalizability from static to dynamic environments. It is the percent relative change between static and dynamic navigation performance. We measure RCS for various $\Lambda$ cases and report scores on SR and SPL. Lower values indicate better generalizability, showing that the agent performance between static and dynamic environments is consistent.
Figure 4: Navigation Generalizability on OTGs (RCS): Relative Change in Success (RCS) measures an agent’s adaptability, comparing dynamic to static TG performance (optimal RCS is $0\%$). Positive values indicate better dynamic performance (chance encounters), while negative values reflect poor adaptability. Experiential-learning methods (RL, LLM+Memory) approach optimal RCS, demonstrating better generalization, while the Oracle and Vanilla LLM perform poorly.
Random $\Lambda$ Semi-Routine $\Lambda$ Fully-Routine $\Lambda$ Avg. OTG
Policy Metric (%) Static OTG RCS (%) Static OTG RCS (%) Static OTG RCS (%) Static OTG RCS (%)
Random Agent SR $66.0$ $74.4$ $+11.2$ $66.0$ $94.7$ $+30.3$ $66.0$ $91.6$ $+27.9$ $66.0$ $86.9$ $+24.1$
TA $33.5$ $24.4$ $-37.3$ $33.5$ $36.6$ $+8.4$ $33.5$ $33.6$ +0.29 $33.5$ $31.2$ $-6.8$
Oracle Agent SR $90.2$ $78.4$ $-13.0$ $90.2$ $75.1$ $-16.8$ $90.2$ $55.9$ $-38.0$ $90.2$ $69.8$ $-22.6$
TA $45.6$ $54.4$ $+16.2$ $45.6$ $50.6$ $+9.8$ $45.6$ $32.7$ $-28.3$ $45.6$ $45.9$ +0.6
RL Agent SR $59.1$ $55.2$ $-6.6$ $59.1$ $57.8$2.3 $59.1$ $55.2$ $-6.7$ $59.1$ $56.0$ $-5.2$
TA $41.6$ $37.3$ $-10.3$ $41.6$ $35.8$ $-14.0$ $41.6$ $32.1$ $-22.9$ $41.6$ $35.1$ $-15.7$
Vanilla LLM Agent SR $28.5$ $43.3$ $+34.1$ $28.5$ $47.2$ $+39.6$ $28.5$ $34.2$ $+16.7$ $28.5$ $41.6$ $+31.4$
TA $24.3$ $30.7$ $+20.9$ $24.3$ $31.3$ $+22.3$ $24.3$ $21.7$ $-10.8$ $24.3$ $27.9$ $+12.9$
LLM + Mem. Agent SR $50.6$ $49.4$2.4 $50.6$ $53.2$ $+4.8$ $50.6$ $47.9$5.4 $50.6$ $50.2$0.9
TA $36.9$ $33.8$8.5 $36.9$ $34.6$6.3 $36.9$ $32.4$ $-12.2$ $36.9$ $33.6$ $-9.0$
Table 3: Navigation on OTGs: We evaluate $5$ popular embodied navigation approaches on various OTGs, and also on a static graph. These values are used to compute RCS using the equation provided in [17]. An RCS value closer to $0\%$ indicates good generalizability to new conditions. Highlighted in green are the best RCS values for SR, while ones in violet are the best values for TA. Observe that the LLM + Memory agent shows the best SR generalizability overall, followed very closely by the RL agent. Both these agents employ experiential learning to learn about object transit patterns and improve their performance.
6.3 Navigation Inference
Table 3 presents the results of our agents on the navigation task on OTGs. We run experiments on $10$ Matterport3D scans, with $10$ randomly selected starting nodes $r\in R$, $20$ episodes, and each episode having $20$ timesteps. There are $4$ key inferences from these observations to understand the generalizability of the various navigation approaches:
Experiential Navigation Approaches Generalize Well: Figure 4 presents a comparison of all approaches on RCS. LLM + Memory and PPO both have a relatively low RCS score, meaning that navigation agents perform comparably in stationary graphs and OTGs. Both methods rely on past observations to improve their decision-making, showing that data-driven approaches can be generalizable well. In contrast, the greedy heuristic Oracle agent and the zero-shot Vanilla LLM lack experience or memory, with the former relying on privileged knowledge and latter utilizing learned commonsense cues for decision-making. This result highlights learned decision-making improves generalizability, and we can infer this to be a key component for the real-world deployment of embodied agents.
Zero-shot Methods Camp to Find Objects: In the Vanilla LLM approach which is zero-shot without memory, we observe that the agent finds more objects in the OTG than in the static environment, as clearly indicated by the higher OTG SR scores across all $\Lambda$s. In Table 3, we see that in the Fully-Routine case for Vanilla-LLM, the RCS for SR is $+16.7\%$ but the RCS for TA is $-10.8\%$. These results suggest that while Vanilla-LLM is finding more objects in OTGs than in static environments, its trajectories show lesser overlap with the ground-truth paths ($\Pi^{*}$). Upon inspection, we observe that this agent has learned to camp at privileged nodes that are in the paths of many object routes. One possible explanation is that the LLM is relying solely on common-sense knowledge and to go to areas where most objects can be found, rather than trying to learning object routes. We can consider camping as an instance of reward hacking, where we would not want an agent to camp at one location in anticipation of encountering a target. This challenge opens up exciting research directions in reward design for navigation on OTGs.
Greedy Heuristic Struggles With Shifting Objects: The navigation approach of the agent heavily relies on the static placement of objects to achieve high performance. This case is best shown in the Oracle agent. At each timestep, the oracle takes a step towards the closest target. However since targets are on the move, a greedy oracle tends to take a suboptimal path. The agent is heading in one particular direction towards a certain target object which was nearest, but then suddenly switches to another direction due to the nearest unseen object changing.
Entropy Influences Performance: In all policies barring the Random agent baseline, we observe that the performance on a Fully Routine $\Lambda$ exhibiting low entropy is usually worse than the Semi-Routine or Random $\Lambda$ cases. This is intuitive, and can be attributed to the agent encountering more targets in environments with higher object transit entropy. Despite this, we observe a strong performance of the RL agent in being entropy-agnostic, with very similar values across all transit cases on SR and SPL.
7 Conclusion, Limitations, and Further Work
We present Object Transition Graphs or OTGs, a generalization of the topological graphs to dynamic environments with shifting objects. Unlike static topological graphs where an agent is expected to navigate to stationary target objects, OTGs are realistic and challenging, with multiple portable objects transiting on a topological graph. Our dynamic task fundamentally challenges the static-scene assumption common in embodied navigation literature.
We first introduce a novel approach to transform topological graphs to OTGs by developing structured object transition scenarios inspired by human habits. Our approach can be used to transform any topological graph into an OTG, and we transform graphs extracted from MP3D for this purpose. On these OTGs, we measure the generalizability of various navigation approaches performing a multi-object finding task. In comparing performance on static TGs and OTGs, we infer the generalizability of these approaches to dynamic settings with non-stationary targets. We find that the type of navigation employed greatly influences agent performance, with experiential approaches that have the means to learn from memory generalizing the best. In one case, we notice an instance of reward hacking via camping behavior, motivating future research in reward design for OTGs. Code and data, including the dynamized MP3D dataset, will be made publicly available to foster research in this exciting new direction.
A limitation of our current setup is that we independently treat visually grounding small target objects as a separate problem. Coupling navigation on OTGs with visual grounding is a future direction. Another interesting direction in this regard is to utilize generative image modeling with OTG navigation on 2D simulators such as Minigrid [13], allowing us to scale up experimentation and training. Finally, our research paves the way for generalizability experiments in the context of dynamic scenes, and future work will look at studying the sim2real transfer of navigation policies trained on OTGs to a real world environment.
Supplementary Material
Supplementary Material
8 Dataset Details
While the agents introduced in Section 2 tackle the temporal dynamism, i.e., identifying where and when the target object is likely to lie at a particular location, visual grounding is a separate task tackling spatial dynamism. An ideal agent performs both these tasks simultaneously, verifying at each timestep if a portable target exists by running a grounding model. As navigation decisions are not directly impacted by the performance of a grounding model, we study temporal and spatial dynamism independently.
The dataset contains temporal and spatial modifications on the Matterport3D scans. Further, we provide code and useful tools for implementing these modifications.
Scan ID # of Nodes # of Edges # of Rooms
QUCTc6BB5sX 145 248 28
8194nk5LbLH 20 32 6
TbHJrupSAjP 114 221 28
2azQ1b91cZZ 215 531 30
oLBMNvg9in8 111 185 31
zsNo4HB9uLZ 53 84 17
EU6Fwq7SyZv 78 166 19
X7HyMhZNoso 84 143 25
x8F5xyUWy9e 43 86 10
Z6MFQCViBuw 58 91 18
Table 4: Summary statistics of scans used for Table 3, including each scan’s ID, number of nodes, edges, and rooms. If a scan has multiple rooms of the same type (e.g. two bathrooms), each instance is counted separately from the total.
For the spatial placement, we provide a dataset of $10,500$ modified images taken across $10$ randomly chosen Matterport3D scans, with $21$ portable objects randomly oriented and positioned $5$ times within each image. Table 4 summarizes the different Matterport3D scans chosen. To emulate realism such that the portable objects are not floating in the sky or on the ceiling, they are placed on a large segmented area below the center of the image. We also provide the bounding boxes pertaining to the portable object in each image.
8.1 Temporal: Implementing Object Placement Strategies
In this section, we give more details on how we implemented the dynamic Matterport3D environment.
Matterport3D Modifications:
Each Matterport3D (MP3D) scan represents a household environment consisting of a set of panoramic view points. Along with the viewpoints (or nodes), we are also given the exact 3D position as well as the relative distance between them. For modifying the MP3D environment, we first construct topological graphs of each scan, with nodes containing the position and the panoramic image, and edges containing relative distance between them. We consider 10 different scans chosen from the REVERIE [36] and R2R [2] unseen validation splits for inference. We choose scans according to these datasets as they contain a variety of rooms for us to populate (Refer Table 1 in the main paper).
The nodes of the topological graphs are then updated at each timestep with the portable objects according to the object placement scenario that has been chosen.
Strategy Overview:
We compute trajectories of the portable objects in an offline manner. For a fixed random seed $s$ and for each portable object $o_{p}\in O_{p}$, we create a sequence of nodes from the graph. In the random transit scenario, we choose any node in the graph. In the routine and semi-routine scenarios, we only choose nodes from plausible rooms. If the node chosen is not the same as the current node, we find the shortest path between the current node and the target node and add the nodes in the path to the sequence representing the trajectory. Once $o_{p}$ gets to the target node, it stays there for either $2$ or $3$ timesteps, determined by a random number generator. For each timestep $o_{p}$ stays at the node, we add the node to the trajectory sequence. After the staying period is over we select a new node and repeat the process until we hit $T$ timesteps.
We take the resulting trajectories and restructure the data to model an evolving graph over a period of $T$ timesteps. The resulting structure is a nested dictionary representing the changing graph. The keys of the outer dictionary are the timesteps [$1,...,T$]. The inner dictionary for each timestep $t$ has the node string ids as the keys and the values are the list of portable objects at that node at $T$. At each timestep $t\in[1,...,T]$ when the agent reaches a node, we simply use $t$ and the node id to retrieve the list of portable objects the agent is currently observing.
8.2 Spatial: Implementing Visual Grounding
In this section, we talk about our strategy for spatial placement of multiple portable target objects on MP3D scenes. After placing portable objects at various locations, we perform visual grounding using three Open-World Object Detection models for inference. We also show further ablations using VLMs.
Generating Spatial Data:
We randomly choose $100$ skybox images taken across the $10$ MP3D scans mentioned in the previous subsection. While selecting these images are not of the ceiling or the floor, and have objects pertaining everyday scenes. We then pick stock images of all the $21$ portable objects present in Table 1 of the main manuscript. These images are shown in Figure 10.
(a) Portable Target: Bowl
(b) Portable Target: Playing Cards
Figure 5: OWL-ViT Detection Failures: Note the poor detection accuracy of OWL-ViT. On both images, it completely misses the object when prompted with a text containing the target label. We attribute these failures to unnatural images being generated by pasting target objects onto the scene.
(a) Portable Target: Playing Cards
(b) Portable Target: Dumbbells
Figure 6: GLIP Detection Accuracy: Note the superior detection accuracy, despite the low coverage on GLIP. On the image on the laft and right, playing cards and dumbbells are detected correctly with a high accuracy.
(a) Portable Target: Screwdriver
(b) Portable Target: Toothbrush
Figure 7: YOLO Detection Failures: Note the poor detection accuracy of YOLO. On the image on the left, the screwdriver is grounded correctly, but the object name predicted is “Cow”. On the image on the right, YOLO completely misses the target object toothbrush, and instead grounds and labels something else.
For each of the skybox images, we first run FastSAM [56] and segment out the largest regions below the center of the image. We then randomly orient the portable object and place it at a random position on one of these segments. Choosing segments below the center of the image is necessary to ensure that the portable objects being placed follow commonsense, and do not hang in the ceiling. This is done $5$ times for each of the $21$ objects on $100$ images, to get $10500$ images in total. While placing the images, we also store the boundaries of the overlay as the ground truth bounding box for computing Intersection over Union (IOU) scores.
Examples of this generated data is shown in Figure 8. Note that the same image can have multiple target segmentations (bed and cushion for example), with different objects placed on them at different orientations.
(a) Dice on Bed
(b) Mug on Cushion
(c) Flashlight on Floor
(d) Charger on Table
Figure 8: Various Spatial Placement in our Dataset: Figures (a) and (b) look at different arrangements of objects in the same room. Figure (c) shows a flashlight on the floor, and figure (d) shows a charger on the table. We cover various orientations and rotations of portable objects being placed on various objects in different MP3D scans.
(a) bowl
(b) dice
(c) dumbbell
(d) flashlight
(e) glasses
(f) hat
(g) salt and pepper shakers
(h) screwdriver
(i) toothbrush
(j) remote
(k) playing cards
(l) phone
(m) phone charger
(n) notebook
(o) mug
(a) laptop
(b) jumprope
(c) headphones
(d) wallet
(e) water bottle
(f) wristwatch
Figure 10: Portable Objects: A list of all the 21 portable objects we consider for our task.
Open-World Visual Object Grounding
We perform Visual Grounding on the spatial placement dataset using various Open-Vocabulary Object Detection Models. For each model, we assess the Detection Coverage, which is the percentage of images where the portable object was found, the Detection Accuracy, the accuracy of the predicted detections, and finally Mean IOU, which gives us the overlap between the ground truth bounding box of the placed target and the predicted box. The results for this are shown in Table 5.
Approach Detection Coverage Detection Accuracy MIOU
OWL-ViT[30] 100% $3.04\%$ $0.352$
GLIP [24] $7.11\%$ 100% 0.489
YOLO-World [12] $32.94$% $0.2$% $0.130$
Table 5: Comparison of Open-World Object Detaction Approaches
We make some interesting observations. OWL-ViT [30] outperforms GLIP [24] and YOLO-World [12] when it comes to Detection Coverage, which means that it consistently detects an object, irrespective of whether it is right or wrong on the image. This is evident since it takes a target object label is given as a prompt to OWL-ViT, so it detects something in the image, even if the confidence is low. Figure 5 shows some examples of this.
GLIP, while having lower detection coverage, has a perfect detection accuracy, meaning that when GLIP does detect something, it usually is correct. The high MIOU score associated with GLIP also corroborates with this fact. Figure 6 shows some example results of this.
We infer the YOLO-World model with custom vocabulary consisting of all the portable object names. Despite this, YOLO still performs the worst among the three, with poor detection accuracy and coverage. Figure 7 showcases one of these results. Notice that in the first image screwdriver is grounded at the right location, but the label is “Cow”. In the second image, both the predicted target (fire hydrant) and the grounding is wrong.
We note that the poor detection as measured by MIOU could be a result of unnatural scenes that are different from those which these models might have been trained on. Beyond this, grounding small, portable targets that could be present on various surfaces in a household environment still remains a challenge. We will be releasing this dataset to foster research in this area.
(a) Phone in bathroom
(b) Headphones on bed
Figure 11: We evaluate the spatial reasoning capabilities of VLMs in grounding small portable targets in houses. The question we ask is “What is the closest number to the $<$INSERT OBJECT$>$ in this image?” and it answers with a number.
8.3 Spatial Reasoning with Vision Language Models (VLMs)
We perform visual grounding using 4 popular Vision Language Models (VLMs): InstructBLIP [28], LLaVA 1.5 [27] and GPT-4o [1]. For InstructBLIP, we experiment with two different language backends: Flan-T5, which has been finetuned on various language tasks including question answering, and Vicuna-13B, which has been trained specifically on conversational data.
Figure 11 illustrates the images we pass to these models to validate their spatial reasoning capability. Given an image containing a small portable target for grounding, we first paste a number grid from $[1-9]$ onto the image. Prior research by Sathyamoorthy et. al [42] and Nasiriany et. al [33] has utilized this technique to guide robots in indoor as well as outdoor scenarios.
In our case, we evaluate a subset of $1000$ images in our dataset containing various objects by asking each image the query - “ Where is the $<$objects$>$ in this image? Reply with the corresponding grid number from [1-9].”, and then evaluate the response we get with the ground truth. Since we create the dataset by pasting target objects onto them, this is readily available to us. Table 6 presents the accuracies for this task the various VLMs.
Model Accuracy (%)
InstructBLIP + Vicuna-7B 15.38
InstructBLIP + Flan-T5-XL 4.32
LLaVA v1.5 + Vicuna 13B 58.62
GPT 4o-mini 75.89
GPT 4o 98.32
Table 6: VQA Model Accuracies: Note the superior performance of GPT 4o on our number grid based spatial reasoning task. The remaining VLMs show inferior performance, with a Vicuna backend performing better than the Flan one; this can be attributed to Vicuna being predominantly trained on conversational data that might contain such queries.
9 Experimental Details
9.1 LGX Inference
We utilize the official implementation of LGX  [^1] https://github.com/vdorbala/LGX and incorporate it into our modified MP3D environment. Briefly, at each timestep, LGX scans the node for objects, and asks an LLM for directions to reach a target. In our case, we seek to maximize finding portable targets. As such, we prompt GPT-4 with the following base prompts -
System Prompt - “I am a smart robot trying to find as many portable objects as I can at home.”
User Prompt - “Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
The $<$OBJECT_LIST$>$ here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX. Additionally, we if a portable object is present at a certain node at a given timestep, we add it to this list.
The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions.
For the memory-enhanced LGX case, we modify the System Prompt, with memory. We additionally ask -
System Prompt:
“I have seen the following objects and taken the following actions so far -
1. $<$OBJECT_LIST$>$: ACTION
2. $<$OBJECT_LIST$>$: ACTION
User Prompt:
“Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
10 Code and Dataset
We provide an anonymous link to our Code: https://anonymous.4open.science/r/otg-1AC0. The dataset will be present as a downloadable link in the code repository.
Furthermore, please see the attached video in the zip file for more information and demonstrations.

### v4

@@
-Right Place, Right Time! Dynamizing Topological Graphs for Embodied Navigation
+TAS: A Transit-Aware Strategy for Embodied Navigation with Non-Stationary Targets
 Abstract
-Embodied Navigation tasks often involve constructing topological graphs of a scene during exploration to facilitate high-level planning and decision-making for execution in continuous environments.
-Prior literature makes the assumption of static graphs with stationary targets, which does not hold in many real-world environments with moving objects.
-To address this, we present a novel formulation generalizing navigation to dynamic environments by introducing structured object transitions to dynamize static topological graphs called Object Transition Graphs (OTGs).
-OTGs simulate portable targets following structured routes inspired by human habits.
-We apply this technique to Matterport3D (MP3D), a popular simulator for evaluating embodied tasks.
-On these dynamized OTGs, we establish a navigation benchmark by evaluating Oracle-based, Reinforcement Learning, and Large Language Model (LLM)-based approaches on a multi-object finding task.
-Further, we quantify agent adaptability, and make key inferences such as agents employing learned decision-making strategies generalize better than those relying on privileged oracle knowledge.
-To the best of our knowledge, ours is the first work to introduce structured temporal dynamism on topological graphs for studying generalist embodied navigation policies.
-The code and dataset for our OTGs will be made publicly available to foster research on embodied navigation in dynamic scenes.
-Figure 1: Object Transition Graph (OTGs): We introduce dynamism to topological graphs via portable targets following structured object transition scenarios.
-In this representative figure, the watch (in red) moves from node N1 at T=10:00 AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM.
-We study the performance of various navigation agents in different object transition scenarios, and establish their generalizability.
+Embodied navigation methods commonly operate in static environments with stationary targets.
+In this work, we present a new algorithm for navigation in dynamic scenarios with non-stationary targets.
+Our novel Transit-Aware Strategy (TAS) enriches embodied navigation policies with object path information.
+TAS improves performance in non-stationary environments by rewarding agents for synchronizing their routes with target routes.
+To evaluate TAS, we further introduce Dynamic Object Maps (DOMs), a dynamic variant of node-attributed topological graphs with structured object transitions.
+DOMs are inspired by human habits to simulate realistic object routes on a graph.
+Our experiments show that on average, TAS improves agent Success Rate (SR) by 21.1 in non-stationary environments, while also generalizing better from static environments by 44.5% when measured by Relative Change in Success (RCS).
+We qualitatively investigate TAS-agent performance on DOMs and draw various inferences to help better model generalist navigation policies.
+To the best of our knowledge, ours is the first work that quantifies the adaptability of embodied navigation methods in non-stationary environments.
+Code and data for our benchmark will be made publicly available.
 1 Introduction
-Embodied navigation tasks often rely on topological graphs (TGs) for decision-making.
-Visual-Language Navigation (VLN) tasks such as REVERIE [36] and R2R [2, 11] utilize TGs extracted from MP3D [8] to follow human guidance.
-Similarly, methods for ObjectNav, PointNav, and ImageNav often employ TGs  [20, 53, 14, 55, 10, 22, 38], where nodes represent locations or landmarks, and edges define traversable paths.
+Embodied navigation involves an agent moving in unseen physical or virtual environments to carry out human-guided tasks such as locating objects [4, 36, 10, 33], following instructions [19, 13, 1, 31], or reaching a point in space [49, 45].
+State-of-the-art methods in this field use Reinforcement Learning (RL)  [47, 45, 20] and Large-Language Model (LLM) based schemes [39, 10] to achieve superior task performance, and these policies also extend to multi-object finding [9, 44].
+Topological graphs (TGs) are a common facet of such techniques and are used to store environmental details for decision making.
+Embodied instruction following tasks such as REVERIE [31]and R2R [1, 13] utilize TGs extracted from MP3D [8] for navigation.
+Similarly, tasks like ObjectNav, PointNav, and ImageNav also employ TGs [20, 46, 14, 48, 11, 22, 34] where nodes represent locations or landmarks, and edges define traversable paths.
 These graphs provide a structured way to plan and execute navigation policies.
-A key limitation of these existing embodied navigation paradigms is their reliance on static graphs, where objects within nodes remain stationary.
-This static assumption is unrealistic, as real-world environments are inherently dynamic, where users frequently move small, portable objects from place to place, such as shifting their phone or wallet between rooms throughout the day.
+A key feature of current embodied navigation paradigms [44, 9, 18, 12] is their reliance on static graphs, where objects within nodes remain stationary.
+This static assumption is limiting, as real-world environments are inherently dynamic, where users frequently move portable objects from place to place over long periods, such as shifting their phone or wallet between rooms throughout the day.
 Studies on robots in collaborative environments  [29, 2, 16] suggest that human object placements follow structured patterns shaped by routines and habits, introducing a degree of entropy in object locations over time.
-Despite this, existing navigation benchmarks fail to model such temporal variability, limiting their applicability for real-world deployment.
-In the ObjectNav task [18, 4] for instance, common state-of-the-art navigation approaches involve Reinforcement Learning (RL) [54, 52, 4, 47] and LLM-based zero-shot methods [18, 45, 6] to find stationary objects.
-RL policies for multi-object settings [50, 9] have also been introduced, but these are also based on static environments.
-For real-world deployment, these approaches must generalize, or maintain performance in dynamic scenes with shifting targets.
-Generalizability in prior embodied work has been described in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to static environments that were unseen or used synthesized guidance language [57, 26, 23, 17].
-However, there has been little to no work that defines generalizability in the context of adapting to dynamic settings with moving objects.
-The evaluation of these prior approaches on dynamic settings with non-stationary objects remains to be seen.
-To address this gap, in this work, we introduce Object-Transition Graphs (OTGs), a dynamic variant of topological graphs that incorporates structured object transitions over time.
+Despite the important applicability of non-stationary environments, existing navigation methods fail to model such temporal variability, limiting their applicability for real-world deployment.
+Figure 1: Transit-Aware Strategy (TAS) We introduce TAS to tackle embodied navigation in non-stationary environments.
+In this figure, an embodied agent is tasked with finding a wallet that changes positions in the environment between 9:$0$ and 9:20.
+$P_{\text{wallet}}$ represents the transit route of the wallet.
+TAS makes agents “transit-aware” by modeling object transit information to augment the agent’s navigation policy for temporally-informed decision making.
+In the ObjectNav task [18, 3], for instance, SotA navigation approaches all involve finding stationary targets using RL [47, 45, 3, 41] and LLM-based zero-shot methods [18, 39, 6].
+RL policies for multi-object settings [44, 9] have also been proposed, but these are also based on static environments.
+In contrast, navigating to non-stationary targets that move as the agent moves necessitates the temporal modeling of object transit to guide the agent.
+For example, OVRL-V2 [47] uses CLIP text embeddings and ViT-processed RGB images, while DD-PPO [45] similarly relies on GPS readings and RGB-D sensors; both use only static observations of the scene.
+Their reward functions similarly assume target stationarity, with OVRL-V2 and DD-PPO penalizing absolute distance changes to the goal.
+Using these approaches directly in non-stationary environments shows poor adaptability because the distance to a target fluctuates as the agent moves, producing a high-variance signal that hinders policy convergence.
+For real-world deployment, these approaches must generalize, i.e., maintain performance in dynamic scenes with shifting targets.
+Prior embodied work describes generalizability in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to improved performance in unseen environments or unique synthesized guidance language [50, 25, 23, 17].
+Little to no work that defines generalizability in the context of adapting to dynamic settings with non-stationary targets.
+Main Results.
+To address these issues, we present a Transit-Aware Strategy (TAS) that informs embodied navigation policies with temporal “awareness” on target positions (Fig.
+1).
+Rather than naively chasing target objects, TAS augmented agents learn to intercept target routes for better convergence.
+To evaluate TAS, we introduce Dynamic Object Maps (DOMs), a dynamic variant of topological graphs incorporating structured object transitions.
 Figure 2 illustrates this concept.
-Unlike traditional TGs, where navigation is performed on a static graph, OTGs introduce a temporal component, where objects shift positions while the underlying graph structure remains constant.
-This forces an agent to synchronize its trajectory with temporal object movements; its adaptability to change allows us to measure its generalist navigation performance.
-In dynamizing static topological graphs, our work seeks to redefine how we evaluate decision-making on embodied agents, by emphasizing not just where to navigate but also when.
-In this work, we define the generalizability of a navigation approach by comparing the performance between static TGs and their OTG counterparts.
-This allows us to quantify an agent’s robustness to dynamic changes, providing us with key insights on the effectiveness of graph-based navigation schemes in real-world settings.
-Main Results.
-We generalize embodied navigation to dynamic environments by 1) introducing structured object transitions to make static topological graphs dynamic, and 2) presenting a benchmark for evaluating the adaptability of state-of-the-art navigation policies in such dynamic settings.
+DOM node attributes change over time, i.e.
+target objects move from node to node at each timestep, while the edges remain fixed.A successful agent must synchronize its path with these object routines, which requires learning and anticipating object movement patterns during exploration.
+This ability reflects an agent’s generalization to non-stationary environments.
+Figure 2: Dynamic Object Maps (DOMs): We introduce dynamism to topological graphs via portable targets following human-habit inspired routes.
+In this representative figure, the watch (in red) moves from node N1 at T=10:$0$ AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM.
+We benchmark navigation agents under different object transition scenarios and present a Transit-Aware Strategy (TAS) to improve their performance.
+Our work seeks to redefine success as defined in traditional embodied navigation tasks, by emphasizing not just on where to navigate but also when to get there i.e.
+right place, right time.
 Our contributions are summarized as follows:
-  • Dynamizing Topological Graphs: We propose a novel framework for introducing structured object transitions on static topological graphs, transforming them into Object Transition Graphs or OTGs.
-Inspired by human habits, our approach defines three distinct object transition scenarios to govern the movement of small, portable objects in a scene, providing varying levels of entropy.
-Our method applies to any static topological graph, including those obtained from Matterport3D (MP3D) [8] and HM3D [37].
-Our dynamized MP3D environments and code will be released as an open-source benchmark to foster research in this area.
-  • OTG Benchmark: We introduce a benchmark for evaluating navigation performance on OTGs, comparing heuristic, RL, and LLM-based agents across structured object transition scenarios on a multi-object finding task.
-Measuring performance of agents on OTGs provides insights into the adaptability of graph-based approaches in dynamic environments with non-stationary objects.
-  • Generalization in Dynamic Environments: We make several key inferences on the generalizability of various navigation approaches by comparing their performance on OTGs against static TGs.
-In particular, we observe that strategies relying on privileged knowledge can lead to suboptimal decision-making in dynamic environments, while approaches that improve agents with experience and collected observations better generalize to OTGs.
+  • Transit-Aware Strategy (TAS): We propose TAS, a novel strategy to enhance the capabilities of embodied agents to find non-stationary targets.
+TAS equips an agent with object transit routes, enabling them to synchronize their paths with moving targets.
+We develop two TAS-enhanced agents TAS-RL and TAS-LLM, that modify existing RL [45] and LLM [18] policies with temporal transit cues.
+  • Dynamic Object Maps (DOMs): We propose an algorithm to introduce structured object transitions on static topological graphs, transforming them into Dynamic Object Maps or DOMs.
+Inspired by human habits, we define three distinct object transition scenarios to govern the movement of portable objects in a scene, providing varying levels of entropy.
+DOMs apply to standard TGs from datasets like Matterport3D (MP3D) [8] and HM3D [32].
+  • Generalizability Benchmark: We evaluate $6$ different navigation agents on a multi-object finding task across static and DOM environments.
+Our TAS-RL and TAS-LLM agents show an average improved Success Rate (SR of $22.9\%$ and generalize better by $49.8\%$ in terms of Relative Change in Success (RCS)[17] over non-TAS counterparts.
 2 Related Works
-Figure 2: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table 1) and transit scenarios (Table 2).
+Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [37, 27, 5] and socially aware navigation [7, 26], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time.
+In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments with stationary targets, with limited research on handling dynamic targets  [13, 38].
+In our work, we introduce dynamism on a graph-level, where agents must reason over long-term spatio-temporal changes of objects for navigation decisions.
+While recent simulators like Habitat 3.0 [30] include object rearrangement tasks, these primarily focus on agents modifying the environment and not finding non-stationary targets on a graph.
+Planning in dynamic environments has been studied in the past [42, 52, 28], with recent approaches even utilizing LLMs in conjunction with multi-arm bandits [15].
+These schemes usually propose memory augmentations with hierarchical planning procedures and frame the problem from an obstacle avoidance standpoint [42, 24, 51].
+In contrast, our work considers the case where the objects themselves are non-stationary.
+Rudra et.
+al.
+[36] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints.
+In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to finding stationary targets at each timestep.
+In contrast, we tackle a truly dynamic case, where objects move even during the episode.
+This definition adds a layer of complexity as the embodied agent must now plan to navigate towards a shifting target, requiring an explicit modeling of the object’s transit.
+Kurenkov et.
+al.
+[21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations.
+They also performed target object-finding experiments, but the target was stationary as the agent moved.
+In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs.
+Wang et.
+al in [43] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments.
+In contrast, our work presents a transit-aware strategy to improve navigation performance in non-stationary environments.
+3 Method: Finding Non-Stationary Targets
+In this section, we first our approach to developing an environment with non-stationary targets with human-habit inspired transit scenarios.
+We then present our strategy to improve embodied navigation performance on these environments by integrating agents with transit-awareness.
+3.1 Defining Transit Scenarios $\Lambda$:
+Human habit formation via object placement can be attributed to cognitive offloading, where individuals rely on the environment to reduce memory demands [35, 40].
+This reliance on the environment influences the spatial distribution of objects.
+Frequently used objects (like keys or phones) might move more flexibly across rooms (high entropy), while function-specific objects (like dumbbells or toothbrush) might be more constrained (low entropy).
+We use this structured transition behavior to define an object’s transition entropy, which we use to formulate $\Lambda$.
+We consider three object transit scenarios at decreasing levels of transit entropy, Random, Semi-Routine, and Fully-Routine.
+These are outlined in Table 1 below.
+Scenario ($\mathbf{\Lambda}$) Fixed Rooms Fixed Paths Entropy
+Static N/A None
+Random High
+Semi-Routine Medium
+Fully-Routine Low
+Table 1: Object Transit Scenarios $\mathbf{\Lambda}$: Portable targets transition in the graph under structured scenarios inspired by human habits.
+In the random case, the portable objects can move to any room at any time during each episode.
+In the routine cases, the rooms that the target objects can travel to are fixed (except during transit).
+The static scenario is a baseline with stationary items i.e., zero entropy.
+Figure 3: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table I in Appendix) and transit scenarios (Table 1).
 Here, a mug is placed on a kitchen node $N_{3}$ at timestep $T=53$, but moves to a bedroom node $N_{5}$ at timestep $T=55$ via node $N_{4}$.
@@
 Multiple objects could also be at the same node (hat and mug in $N4$).
-Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [41, 31, 5] and socially aware navigation [7, 29], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time.
-In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments, with limited research on handling dynamic targets  [11, 44].
-With OTGs, we introduce a form of structured temporality in dynamic environments where agents must reason over spatio-temporal changes, extending high-level planning to dynamic scenes with shifting targets.
-While recent simulators like Habitat 3.0 [35] include object rearrangement tasks, these primarily focus on agents modifying the environment rather than adapting to dynamic object placements over time.
-Planning in dynamic environments has been studied in the past [48, 59, 32], with recent approaches even utilizing LLMs in conjunction with multi-arm bandits [15].
-These schemes usually propose memory augmentations with hierarchical planning procedures and frame the problem from an obstacle avoidance standpoint [48, 25, 58].
-In contrast, our work considers the case where the objects themselves are non-stationary.
-Rudra et al.
-[40] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints.
-In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to an object-finding navigation task in a static environment.
-In contrast, we tackle a truly dynamic case, where objects are moving even during the episode.
-This definition adds a layer of complexity as the embodied agent must now navigate towards a constantly shifting target object, for which it needs to identify specific routines and object movement patterns in the environment.
-Kurenkov et al.
-[21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations.
-They also performed target object-finding experiments, but the environment was static as the agent moved.
-In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs.
-Furthermore, Wang et al.
-[49] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments.
-In contrast, our work focuses on converting topological graphs into object transition graphs, and comments on the generalizability of embodied navigation approaches in dynamic environments.
-3 Object Transition Graph (OTG):
-We define Object Transition Graphs (OTGs) is an evolutionary undirected graph where the underlying graph structure remains constant, but the objects present in its nodes evolve over time.
+3.2 Dynamic Object Maps (DOMs):
+We define Dynamic Object Maps (DOMs) as node-attributed dynamic graphs where edge values remain constant, but the node attributes (objects) evolve over time.
+DOMs extend attributed graphs with temporal object dynamics.
 Formally, let $G=(V,E)$ be a topological graph structure with nodes $V$ representing spatial locations (rooms and objects), and edges $E$ representing traversable paths between the nodes.
@@
 Then, at each timestep $t$, $v\in V$ is associated with a set of portable target objects $O_{v}(t)\subseteq\mathcal{O}$.
-While the graph topology remains static, the object assignments $O_{v}(t)$ evolve via a time-dependent transitioning style $\Lambda$, to give us an object transition function $O$ as,
+While the graph edges $E$ remain fixed over time, the node attributes or object assignments $O_{v}(t)$ evolve via a time-dependent transit scenario $\Lambda$, to give us an object transition function $O$ as,
 $O:V\times\Lambda\rightarrow 2^{\mathcal{O}}$ (1)
 where $2^{\mathcal{O}}$ is a power set representing all possible portable object combinations.
-At each timestep $t$, $O(v,t)$ gives us the portable targets present in each node $v$ based on the transition style $\Lambda$.
+At each timestep $t$, $O(v,t)$ gives us the portable targets present in each node $v$ based on the transit scenario $\Lambda$.
+Figure 3 illustrates this evolving graph setup.
+The initial graph contains portable objects placed according to Table I (See Appendix).
+We then generate object paths $\zeta$ by moving them in the graph at various timesteps for each transit scenario $\Lambda$.
+Algorithm I in the Appendix describes our approach to generating $\zeta$.
+When generating the evolving graph, a pseudo-random seed $s$ helps differentiate the movement of portable objects per episode. Letting $e$ be the episode index,
+[…]
+Note that for the fully routine case, $\zeta$ remains constant for all episodes of the experiment, meaning that the object transition routes remain fixed.
+In the other two cases, $\zeta$ varies per episode according to the room and interval variability described.
+3.3 Transit-Aware Strategy (TAS)
+Many prior embodied navigation tasks work under the premise of “static and stationary”, where targets are assumed to stay at one location throughout the episode.
+This premise is also true with zero-shot LLM-based approaches [39, 18] As such, they are trained with a reward to penalize the agent for being too far from the target [45, 47].
+This can be written as -
+$\mathcal{R}_{\text{dist}}=-\alpha\parallel{v_{a}-v_{o}}\parallel.$
+This standard formulation fails with DOMs, since $v_{o}$ changes during the episode.
+As a result, even though $R_{\text{dist}}$ is a dense reward, the reward signal fluctuates too much to enable meaningful learning.
+This fluctuation is reflected in the poor performance of DD-PPO [45] on DOMs and is discussed later in the results section.
+To improve performance, agents need to be made “aware” of target objects in transit, and this involves integrating them with temporal knowledge of the scene.
+We present modifications to an RL agent and an LLM agent as follows:
+TAS-RL Agent: To make an RL agent transit-aware, we make the following modifications to its setup:
+Observations and Actions: Given a DOM $G=(V,E)$, we provide the agent with a set of paths $\mathcal{P}$ denoting potential object transits.
+$P_{o}\in\mathcal{P}$ is a list of timed vertices $v_{1},\dots v_{n}]$ denoting the position of object $o$ between timesteps $1\dots n$.
+Formally,
+$\mathcal{P}=\left\{P_{o}\;\middle|\;o\in\mathcal{O}\right\},\quad\text{where}\quad P_{o}=\left\{(t_{i},v_{i})\right\}_{i=1}^{n},\;t_{i}\in T,\;v_{i}\in V.$
+When $\mathcal{P}$ is not readily available, it can be initialized with known priors about human habits in the environment.
+In a practical example, if a human moves tends to move their wallet between the kitchen and bedroom between 99:$00$ and 99:2020, the wallet path could be approximated as $P_{wallet}=[v_{9:00},v_{9:01}\dots v_{9:20}]$, where $v_{t}$ indicates the wallet’s position a time $t$.
+The transit scenario $\Lambda$ categorizes $P_{o}$’s temporal structure.
+We also provide the current timestep $t\in T$ and the current position of the agent $v_{c}$.
+Given this information, the agent chooses the next node to go to.
+We employ a maskable action space to ensure that the agent moves only to the neighboring nodes of $v_{c}$.
+Reward Objective: As described earlier, a distance reward $R_{dist}$ provides a poor and highly variable training signal.
+We replace it with a sparse interception reward $\mathcal{R}_{I}$ that is triggered whenever the agent intersects with the target’s path $P_{o}$.
+$\mathcal{R}_{I}=\begin{cases}r_{i},\>\text{if}\>v\in P_{o}\\ 0,\>\text{otherwise}.\end{cases}$
+This reward incentivizes the agent to synchronize its path with that of the target, rather than greedily “chasing” a target to minimize distance.
+$\mathcal{R}_{I}$ along with a standard target finding reward $\mathcal{R}_{f}$ allows for more informed agent pathing for finding non-stationary targets.
+TAS-LLM Agent: LLM-based navigation agents use commonsense reasoning about observations to predict navigable actions.
+In LGX [18] for instance, if the target is a ‘leaf-shaped bowl’, the agent uses observations of regular objects like ‘table’ and ‘sofa’, to determine a direction that is more likely to have the target (the ‘table’).
+In DOMs, object transit scenarios $\Lambda$ similarly use commonsense locations defined in Table I (See Appendix).
+As such, passing previously seen regular objects to the LLM would help the agent make more transit-aware decisions.
+Since LLM agents are predominantly zero-shot and work directly during inference time with no training, we modify the system prompt to make them transit-aware.
+At each timestep $ts_{i}$, we gather the action sequence $a_{0->ts_{i}}$, and a set of objects in the scene (may also include the target) $o_{t}$.
+In the subsequent timesteps, we pass this transit data $\mathcal{T}_{i}=[ts_{i},a_{0->ts_{i}},o_{t}]$ into the system prompt.
+A horizon is maintained to avoid token overflow, where the earliest appended observation is removed at the horizon.
+For both TAS-RL and TAS-LLM, note that the Transit Scenario $\Lambda$ plays an important role in determining agent success.
+An agent finding an object in random transit $\Lambda_{\text{random}}$ would find it a lot harder than one in the routine transit cases $\Lambda_{\text{semi-routine}}$ or $\Lambda_{\text{routine}}$.
+For the former, even if an agent intersects with $\Lambda_{\text{random}}$ and follows the trajectory, it may never find the target is $\Lambda_{\text{random}}$ is spatio-temporally random.
+4 Task: Multi-Object Finding
+To evaluate the benefits of TAS, we focus on the task of finding multiple portable target objects over a fixed timespan.
+This setting adds realism to standard multi-object finding tasks such as Multi-ON [44] and GOAT [9] where targets are stationary.
+In our dynamic setting, we argue that a multi-object finding task is preferable over finding a single portable object for two reasons:-
+  • Realistic Setting: In real-world homes, users frequently move multiple objects simultaneously, leading to dynamically changing object locations [2, 29].
+Assistive agents must track and infer object movements over time to provide meaningful assistance when asked.
+This requires monitoring multiple portable objects (e.g., phones, watches) rather than treating each search as an isolated event.
+  • Lifelong Navigation: A multi-object setup enables an agent to finetune its search policy in real-time, by using knowledge gathered from finding previous objects to better find the next.
+An agent deployed at homes will likely have to face similar conditions, with a user asking it to find multiple targets without ‘resetting’ its position each episode as the case with single-target tasks, much akin to lifelong navigation[46].
+First, we mathematically formalize what navigating on a DOM entails.
+Let $\Lambda$ describe the transitory motions of objects.
+Then, let $O$ be the set of portable objects found over $T$ timesteps for an agent starting at node $r$.
+Then $S=[r,T,\Lambda]$ represents a set of our experimental variables for this task.
+We then define our navigation objective as “Finding the maximum possible number of portable objects $O$ while traversing a dynamic environment represented by SS.”.
+For downstream applications like embodied agents personalizing to households, we believe finding multiple targets to be a good task to study dynamic adaptability, due to dense reward signals [9].
+The setting is also realistic as households have multiple people moving objects around at the same time.
+We study the performance of different navigation policies with varying simulator conditions $S$.
+In each episode, the agent is tasked with finding as many portable target objects as possible by making decisions on a dynamic DOM.
+Let $\tau$ be a finite trajectory representing a sequential list of visited nodes on the DOM, $G$, and let $\tau(t)$ be the $t$-th node in $\tau$.
+We can now define $O(\tau,t)$ as the set of portable objects found at the $t$-th node of $\tau$ at time $t$.
+Given that the objects move throughout the scene, the $t$-th node may have a different set of portable objects present at a different timestep.
+Given a trajectory $\tau$ of length $T$, we can formulate the total set of portable objects found along $\tau$ as $O_{\tau}=\cup_{t=1}^{T}O(\tau,t)$.
+A navigation agent parameterized by policy $\pi$ explores the dynamic DOM by generating a trajectory $\tau_{\pi}$, where each timestep $t$ corresponds to a selected node: $\tau_{\pi}(t)\in V$.
+Specifically, $\tau_{\pi}(t)$ is the node chosen at time $t$ by policy $\pi$.
+We now can write the policy optimization problem as finding an optimal policy $\pi^{*}$ defined as $\pi^{*}=\operatorname*{arg\,max}_{\pi}|O_{\tau_{\pi}}|$.
+Consequently, $O_{\tau_{\pi^{*}}}$ is the maximum number of portable targets that can be found by taking the optimal trajectory.
+5 Experiments and Results
+We perform multi-object finding on the $4$ different $\Lambda$ scenarios defined in Table 1.
+DD-PPO [45] and LGX [18] are baseline RL and LLM agents and are TAS enhanced using the strategy described in Section 3.3 to produce TAS-RL and TAS-LLM agents.
+We also benchmark performance with Random and Oracle agents.
+5.1 DOM Setup
+For our experiments, we convert static topological graphs from the Matterport3D (MP3D) [8] dataset to obtain DOMs.
+Each node contains panoramic images representing the scene, and edges represent the distance between them.
+We determine $4$ hyperparameters that influence performance i.e., $H=[r,E,T,\Lambda]$, where $r\in G$ is a random starting node chosen from $G$, $E$ is is the number of episodes per trial, $T$ is the number of timesteps per episode and $\Lambda$ is the transit scenario.
+Thus, each agent is subjected to a total of $Stationary(1)+Transit(3)]\times r\times E\times T$ experiments.
+We define each trial as a set of episodes over a fixed time from a starting node $r$.
+The maximum possible number of trials for a scan is thus equal to the size of the graph $G$.
+In our subset of $10$ MP3D scans (Ref.
+Appendix for details), $|r|$ ranges from $[20,215]$.
+For each scan, we set $|r|=10$, $E=20$, and $T=30$ in our experiments.
+Although different objects can have different transition scenarios, we set all objects in our experiments to have the same $\Lambda$.
+This allows us to better analyze the generalizability of a navigation scheme and avoids complexities involved with defining per-object $\Lambda$ associations.
+Ground Truth Paths ($\Pi^{*}$): Given a starting node $r\in G$, for each episode $e\in E$, there exist multiple ground truth paths $\Pi^{*}$ over $T$ that the agent could take for collecting the most number of portable targets.
+We compute these paths by brute force, simulating all possible trajectories from $r$ over $T$ timesteps.
+$\Pi^{*}$ is used in our evaluation metrics.
+5.2 Evaluation Metrics
+Success Rate (SR): For each episode, we measure the object-finding performance of an agent by dividing the number of objects found in that episode by the maximum possible number of objects that could have been found with a ground-truth path $\pi^{*}\in\Pi^{*}$: $SR=|O_{\pi}|/|O_{\pi^{*}}|$.
+This metric is similar to the Progress metric in [44].
+In our task, the ground-truth trajectory may not necessarily find all the objects in the given timeframe.
+Trajectory Alignment (TA): To measure path efficiency, we define Trajectory Alignment as the average overlap between ground-truth trajectory $\pi^{*}$ and trajectory taken by the agent across episodes.
+We take the maximum across $\Pi^{*}$: $TA=\min_{\pi^{*}\in\Pi^{*}}\frac{1}{T}\sum_{t\in T}\mathbf{1}_{\pi(t)=\pi^{*}(t)}$, where $\mathbf{1}$ is an indicator function.
+Relative Change in Success (RCS): We use this value taken from [17] to measure the generalizability from static to dynamic environments.
+It is the percent relative change between static and dynamic navigation performance.
+We measure RCS for various $\Lambda$ cases and report scores on SR and TA.
+Lower values indicate better generalizability, showing consistency in agent performance between static and dynamic environments.
+Table 2: Navigation on DOMs: We evaluate 66 embodied navigation approaches on various DOMs to establish a benchmark.
+These values are used to compute RCS using the equation provided in [17].
+An RCS value closer to $0\%$ indicates good generalizability to new conditions.
+Highlighted in green are the best RCS values for SR, while ones in violet are the best values for TA.
+Observe the improved RCS scores of TAS-RL and TAS-LLM agents, indicating better generalizability to non-stationary targets.
+Particularly note that the RL agent performs well in the static scenario, with a significant performance drop when subjected to DOMs.
+The LLM agent shows improved performance on DOMs, but we attribute this to multiple chance encounters with moving targets.
+$\Lambda$ Random $\Lambda$ Semi-Routine $\Lambda$ Fully-Routine $\Lambda$ Average Policy Metric (%) Static DOM RCS Static DOM RCS Static DOM RCS Static DOM RCS Random SR $42.0$ $74.4$ +43.5+43.5 $42.0$ $54.7$ +23.2+23.2 $42.0$ $48.3$ +13.0+13.0 $42.0$ $59.1$ +29.0+29.0 TA $33.5$ $24.4$ 37.3-37.3 $33.5$ $36.6$ +8.4+8.4 $33.5$ $33.6$ +0.29 $33.5$ $31.2$ 6.8-6.8 Oracle SR $90.2$ $78.4$ 13.0-13.0 $90.2$ $75.1$ 16.8-16.8 $90.2$ $55.9$ 38.0-38.0 $90.2$ $69.8$ 22.6-22.6 TA $45.6$ $54.4$ +16.2+16.2 $45.6$ $50.6$ +9.8+9.8 $45.6$ $32.7$ 28.3-28.3 $45.6$ $45.9$ +0.6 RL [45] SR $61.8$ $33.4$ 45.9-45.9 $61.8$ $21.8$ 64.7-64.7 $61.8$ $12.2$ 80.1-80.1 $61.8$ $22.4$ 63.7-63.7 TA $59.8$ $36.5$ 38.9-38.9 $59.8$ $29.8$ 50.1-50.1 $59.8$ $24.1$ 59.6-59.6 $59.8$ $30.1$ 49.6-49.6 TAS-RL SR $59.1$ $55.2$ 6.6-6.6 $59.1$ $57.8$ -2.3 $59.1$ $55.2$ 6.7-6.7 $59.1$ $56.0$ 5.2-5.2 TA $41.6$ $37.3$ 10.3-10.3 $41.6$ $35.8$ 14.0-14.0 $41.6$ $32.1$ 22.9-22.9 $41.6$ $35.1$ 15.7-15.7 LLM [18] SR $28.5$ $43.3$ +34.1+34.1 $28.5$ $47.2$ +39.6+39.6 $28.5$ $34.2$ +16.7+16.7 $28.5$ $41.6$ +31.4+31.4 TA $24.3$ $30.7$ +20.9+20.9 $24.3$ $31.3$ +22.3+22.3 $24.3$ $21.7$ 10.8-10.8 $24.3$ $27.9$ +12.9+12.9 TAS-LLM SR $50.6$ $49.4$ -2.4 $50.6$ $53.2$ +4.8+4.8 $50.6$ $47.9$ -5.4 $50.6$ $50.2$ -0.9 TA $36.9$ $33.8$ -8.5 $36.9$ $34.6$ -6.3 $36.9$ $32.4$ 12.2-12.2 $36.9$ $33.6$ 9.0-9.0
+5.3 Navigation Inference
+Figure 4: Navigation Generalizability on DOMs (RCS): Relative Change in Success (RCS) [17] measures an agent’s adaptability, comparing dynamic to static TG performance (optimal RCS is $0\%$).
+Positive values indicate better dynamic performance (chance encounters), while negative values reflect poor adaptability.
+Both our TAS-enhanced agents approach optimal RCS, demonstrating better generalization to non-stationary environments.
+Table 2 presents the results of our agents on the navigation task on DOMs.
+We run experiments on $10$ Matterport3D scans, each with $10$ randomly selected starting nodes $r\in G$, $20$ episodes, and each episode having $20$ timesteps.
+We draw $4$ key inferences from these observations to understand the generalizability of the various navigation approaches:
+TAS improves Generalizability: Figure 4 presents a comparison of all approaches on RCS.
+TAS-LLM and TAS-RL both have a relatively low RCS score, meaning that navigation agents perform comparably in stationary graphs and DOMs.
+This highlights the effectiveness of incorporating transit-awareness in non-stationary targets.
+In contrast, the greedy heuristic Oracle agent and the zero-shot LLM lack transit information, with the former relying on privileged knowledge and latter utilizing commonsense cues about static observations for decision-making.
+We can infer the impact of TAS as a key component for the real-world deployment of embodied agents.
+Future work will explore the heuristic estimation of human habits as a prior for estimating transit paths for TAS.
+[…]
+Greedy Oracle Struggles With Shifting Objects: The Oracle agent uses a greedy heuristic to select a node towards the closest target object, and this works well in the static scenario.
+On DOMs however, targets are on the move and the greedy oracle tends to take a suboptimal path.
+The oracle heads in one particular direction towards a certain target object which was nearest, but keeps switching directions when another target appears to be closer.
+This behavior is a direct approximation of the distance reward with DD-PPO that causes target chasing and highlights the need for transit awareness.
+Entropy Influences Performance: In all policies barring the Random agent baseline, we observe that the performance on a Fully Routine $\Lambda$ exhibiting low entropy is usually worse than the Semi-Routine or Random $\Lambda$ cases.
+This can be intuitively attributed to the agent encountering more targets in environments with higher object transit entropy (Fig.
+3 in the Appendix).
+Despite this, we observe a strong performance of the TAS-RL agent in this case, with very similar values across all transit cases on SR and TA.
+This indicates that TAS can enhance embodied RL agents to become agnostic to transit scenarios where all target objects follow the same $\Lambda$.
+Future work will study the case of mixed $\Lambda$ with different objects following a different transit scenario.
+6 Conclusion, Limitations, and Further Work
+We present a novel approach to perform embodied navigation in dynamic environments with non-stationary targets.
+Our method enhances RL and LLM-based policies traditionally modeled for stationary targets with a Transit-Aware Strategy (TAS) that enables them to capture object transit information.
+To evaluate our TAS-enhanced agents, we introduce DOMs, or Dynamic Object Maps, an algorithm to translate static topological graphs into dynamic environments with shifting objects.
+Object transitions in DOMs aim to simulate human object placement habits.
+Our work breaks the underlying “static graph, stationary target” assumption common with embodied navigation literature, and presents a navigation benchmark in a dynamic environment with non-stationary targets.
+Our TAS-enhanced agents outperform non-TAS counterparts by $21.1\%$ when measured by average Success Rate in DOMs, while also drastically generalizing better to non-stationary environments by over $44.5\%$ when measured by Relative Change in Success.
+Code and data for our work will be made publicly available, and we hope to foster research in this domain.
+Future work will look at the sim2real transfer of TAS agents to real-world environments.
+Supplementary Material
+7 DOM Algorithm
+To “DOMify” our environments, we use the following algorithm.
+Given a set of portable objects $\mathcal{O}$ and a topological graph $G$, we generate an evolving node-attributed graph trajectory $\zeta$ over $T$ timesteps that follows an object transit scenario $\Lambda$.
+Objects $\mathcal{O}$ and $\Lambda$ use Table 3 as reference.
+We manually populate this table with commonsense knowledge on general object locations.
+This grounded placement plays a key role later in influencing LLM agent performance, since they rely on LLM commonsense knowledge to make navigation decisions.
+Algorithm 1 DOM: Modify $G$ With Object Transitions $\Lambda$
+Input: Topological Graph $G=(V,E)$, Portable Object Set $\mathcal{O}$, Episode Length $T$, Initial Agent Node $r$, Object Transit Scenario $\Lambda$
+Output: Evolving Graph Trajectory $\zeta$
+Initialize: Add $\mathcal{O}$ portable objects to $G_{0}$.
+1:for $t\in[1,\dots,T]$ do
+2:  $G_{t}=G_{t-1}$
+3:  for each object $o\in\mathcal{O}$ do
+4:   Using Equation 1: $O(v,t)\leftarrow O(v,\Lambda)$
+5:   Sample $v_{k}\in V$ from $O(v,t)$ {Following $\Lambda$}
+6:   if $v_{k}\neq v_{o}$ then
+7:    if $\Lambda=\text{Fully-Routine}$ then
+8:     Use fixed trajectory $\delta(t)$
+9:    else
+10:     Compute random path $\delta(t)$ from $v_{o}$ to $v_{k}$
+11:    end if
+12:    Move $o$ along $\delta$ to $v_{k}$ over $t$ timesteps
+13:   end if
+14:  end for
+15:  Add $G_{t}$ to $\zeta$
+16:end for
+17:return $\zeta$
+Table 3: Rooms and Portable Objects: We map 2121 portable objects to a set of household rooms.
+This mapping is used to set destinations for object transit.
+During each episode, objects are placed in various rooms for a range of timesteps.
+Commonly moved objects such as phone, headphones are associated with $9$ different rooms, while less commonly shifted ones such as dumbbells appear only in the Gym.
 Room Portable Objects
@@
 Recreation Playing Cards, Dice, Water Bottle, Headphones, Hat
-Table 1: Rooms and Portable Objects: We map $21$ portable objects to a set of household rooms.
-This mapping is used to set destinations for object transit.
-During each episode, objects are placed in various rooms for a range of timesteps.
-Commonly moved objects such as phone, headphones are associated with $9$ different rooms, while less commonly shifted ones such as dumbbells appear only in the Gym.
-Algorithm 1 Modify $G$ With Object Transitions $\Lambda$
-Input: Topological Graph $G=(V,E)$, Portable Object Set $\mathcal{O}$, Episode Length $T$, Initial Agent Node $r$, Object Density $d$, Object Transit Scenario $\Lambda$, Local Graph Size $q$.
-Output: Evolving Graph Trajectory $\zeta$
-Initialize: Add $\mathcal{O}$ portable objects to $G_{0}$. Calculate the object density $d$ of the local graph $g$ around the initial node $r$. $g$ is all nodes within $q$ steps of $r$.
-Sample $N\in\mathcal{O}$ duplicate new objects and add them to random nodes in $g$ to match required density $d$. $\mathcal{O}^{\prime}=\mathcal{O}+N$ is the total set of objects at $G_{0}$
-1:  for $t\in[1,\dots,T]$ do
-2:     $G_{t}=G_{t-1}$
-3:     for each object $o\in\mathcal{O}^{\prime}$ do
-4:        Using Equation 1: $O(v,t)\leftarrow O(v,\Lambda)$
-5:        Sample $v_{k}\in V$ from $O(v,t)$Following $\Lambda$}
-6:        if $v_{k}\neq v_{o}$ then
-[…]
-8:              Use fixed trajectory $\delta(t)$
-9:           else
-10:              Compute random path $\delta(t)$ from $v_{o}$ to $v_{k}$
-11:           end if
-12:           Move $o$ along $\delta$ to $v_{k}$ over $t$ timesteps
-13:        end if
-14:     end for
-15:     Add $G_{t}$ to $\zeta$
-16:  end for
-17:  return $\zeta$
-3.1 Defining Transition Scenarios $\Lambda$:
-Scenario ($\mathbf{\Lambda}$) Fixed Rooms Fixed Paths Entropy
-Static N/A None
-Random High
-Semi-Routine Medium
-Fully-Routine Low
-Table 2: Object Transit Scenarios $\mathbf{\Lambda}$: Portable targets transition in the graph under structured scenarios inspired by human habits.
-In the random case, the portable objects can move to any room at any time during each episode.
-In the routine cases, the rooms that the target objects can travel to are fixed (except during transit).
-The static scenario is a baseline with stationary items i.e., zero entropy.
-Human habit formation via object placement can be attributed to cognitive offloading, where individuals rely on the environment to reduce memory demands [39, 46].
-This reliance on the environment influences the spatial distribution of objects.
-Frequently used objects (like keys or phones) might move more flexibly across rooms (high entropy), while function-specific objects (like dumbbells or toothbrush) might be more constrained (low entropy).
-We use this structured transition behavior to define an object’s transition entropy, which we use to formulate $\Lambda$.
-We consider three object transit scenarios at decreasing levels of transit entropy:- 1) Random, Semi-Routine, and Fully-Routine.
-These are outlined in Table 2.
-(a) Random Scenario
-(b) Semi-Routine Scenario
-(c) Fully-Routine Scenario
-Figure 3: Object Frequency Visualization: A visual representation of portable target frequencies over several timesteps in various transit scenarios $\Lambda$.
-We simulate portable objects transiting on a topological graph from MP3D.
-In the random case, object frequencies are more scattered (high entropy); in semi-routine, target rooms are fixed, but paths are flexible (medium entropy); and in the routine case, both rooms and paths are fixed (lowest entropy).
-Note the distribution of object frequencies visually correlates to their entropy.
-Algorithm 1 describes our approach for generating an evolving graph trajectory $\zeta$ to dynamize any topological graph $G$.
-Figure 2 further illustrates this setup.
-The initial graph contains portable objects placed according to Table 1.
-We then generate $\zeta$ by moving objects in the graph at various timesteps for each transit scenario $\Lambda$.
-$\zeta$ is precomputed for each transit scenario and stored for later analysis with navigation algorithms.
-When generating the evolving graph, a pseudo-random seed $s$ helps differentiate the movement of portable objects per episode. Letting $e$ be the episode index,
-$s=\begin{cases}1\>\text{(or fixed)}&\text{if movement is fully routine},\\ e&\text{if movement is random or semi-routine}.\end{cases}$ (2)
-Note that for the fully routine case, $\zeta$ remains constant for all episodes of the experiment, meaning that the object transition routes remain fixed.
-In the other two cases, $\zeta$ varies per episode according to the room and interval variability described.
-Figure 3 presents a visual representation of the object frequencies in all three scenarios.
-For our experiments, we dynamize topological graphs obtained from the Matterport3D (MP3D) [8] dataset using Algorithm 1 to obtain OTGs.
-Each node contains panoramic images representing the scene, and edges represent the distance between them.
-Additionally, we have the room type for each node and use them as a template to define the set of portable objects listed in Table 1.
-Note on Visual Spatial Placement.
-Beyond object transitions, we also investigate spatial transformations in placing the objects on the scene.
-Portable objects are small and can be pasted on the scene in various ways.
-We first use FastSAM [56] to obtain segmentation masks for both the scene and the portable target object.
-On the scene segmentations, we filter out a set of $K$ largest segments below the center of the image, to avoid placing floating targets.
-We then place the target objects on a randomly chosen segment $k\in K$ this set.
-This simple placement technique for small objects allows us to modify images of pre-existing topological graphs and makes the visual grounding task challenging enough for comparison against various backgrounds.
-While 3D object placement would be more accurate, this is more cumbersome to setup, and is tangential to the focus of this paper, which is to introduce OTGs to study navigation generalizability.
-More details on our visual grounding experiments with VLMs can be found in the Appendix.
-We will be releasing this dataset of small objects placed on MP3D to foster research in embodied visual grounding.
-4 Task: Multi-Object Finding
-To evaluate how well different navigation methods perform on dynamic OTGs, we aim to assess their generalizability.
-Given a topological graph, the generalizability is the difference in performance achieved by a navigation scheme between when then graph is static and when the graph is dynamic.
-For example in Section 6, we measure the generalizability of PPO-based agents by training an agent in a static environment and training another in a dynamic environment and compare the difference in performance.
-To enable this study, we focus on the task of finding multiple portable target objects over a fixed timespan.
-This setting adds realism to other multi-object finding tasks such as Multi-ON [50] and GOAT [9] by introducing non-stationary targets.
-For dynamic settings such as ours, we argue that a multi-object finding task is preferable over a standard single-object finding setup for two reasons:-
-  • Realistic Setting: In real-world homes, users frequently move objects, leading to dynamically changing object locations [3, 34].
-Assistive agents must track and infer object movements over time to provide meaningful assistance.
-This requires monitoring multiple portable objects (e.g., phones, watches) rather than treating each search as an isolated event.
-  • Lifelong Navigation: A multi-object setup enables an agent to finetune its search policy in real-time, by using knowledge gathered from finding previous objects to better find the next.
-An agent deployed at homes will likely have to face similar conditions, with a user asking it to find multiple targets without ‘resetting’ its position, much akin to lifelong navigation [53].
-First, we mathematically formalize what navigating on an OTG entails.
-Let $\Lambda$ describe the transitory motions of objects.
-Then, let $O$ be the set of portable objects found over $T$ timesteps for an agent starting at node $r$ with an object distribution density $d$.
-Then $S=[r,d,T,\Lambda]$ represents a set of our experimental variables for this task.
-We then define our navigation objective as “Finding the maximum number of portable objects $O$ while traversing a dynamic environment represented by $S$.”.
-For downstream applications like embodied agents personalizing to households, we believe finding multiple targets to be a good task to study dynamic adaptability, due to dense reward signals [9].
-The setting is also realistic as households have multiple people moving objects around at the same time.
-We study the performance of different navigation policies with varying simulator conditions $S$.
-In each episode, the agent is tasked with finding as many portable target objects as possible by making decisions on a dynamic OTG.
-Let $\tau$ be a finite trajectory representing a sequential list of visited nodes on the OTG, $G$, and let $\tau(t)$ be the $t$-th node in $\tau$.
-We can now define $O(\tau,t)$ as the set of portable objects found at the $t$-th node of $\tau$ at time $t$.
-Given that the objects move throughout the scene, the $t$-th node may have a different set of portable objects present at a different timestep.
-Given a trajectory $\tau$ of length $T$, we can formulate the total set of portable objects found along $\tau$ as $O_{\tau}=\cup_{t=1}^{T}O(\tau,t)$.
-A navigation agent parameterized by policy $\pi$ explores the dynamic OTG by generating a trajectory $\tau_{\pi}$, where each timestep $t$ corresponds to a selected node: $\tau_{\pi}(t)\in V$.
-Specifically, $\tau_{\pi}(t)$ is the node chosen at time $t$ by policy $\pi$.
-We now can write the policy optimization problem as finding an optimal policy $\pi^{*}$ defined as $\pi^{*}=\operatorname*{arg\,max}_{\pi}|O_{\tau_{\pi}}|$.
-Consequently, $O_{\tau_{\pi^{*}}}$ is the maximum number of portable targets that can be found by taking the optimal trajectory.
-In the following section, we describe various navigation schemes that we employ to perform the multi-object finding task on OTGs.
-5 Navigation Approaches
-To infer the generalizability of prior navigation models, we emulate a static topological graph environment by keeping portable objects stationary in one location throughout the experiment.
-We then evaluate both the static topological graph environment and the dynamic OTG environment with $5$ relevant navigation agents.
-Random Agent. For a completely random navigation baseline, at each timestep $t$, the agent chooses one of its neighboring nodes to go to for the next iteration.
-Oracle Agent.
-We utilize a greedy heuristic on an agent as an oracle baseline: At each timestep $t\in[1,T]$, we perform a Breadth-First Search (BFS) to find the closest node with an unseen portable object.
-Using the shortest path between the unseen object and its current position, the agent then moves its neighboring node on the given shortest path.
-This approach relies on the agent having access to an “oracle” that tells it about the closest node containing a unique portable target at a given timestep.
-While oracle knowledge is useful since the position of the objects in the environment changes with time, note that there is a possibility that it will not find an object after navigating to an oracle-guided node.
-Reinforcement Learning Agent.
-We use a reinforcement learning-based method, specifically Proximal Policy Optimization (PPO) [43] to define an agent.
-PPO has been used in several previous works in embodied navigation tasks, such as ObjectNav [54, 60, 51] with SOTA results.
-Observation Space: Given the limited prior literature dealing with PPO on a dynamic topological graph, we consider constructing our agent with the following observation space as input to the navigation agent.
-To define the observation space, let $|G|$ be the number of nodes on $G$, and each node is given a unique ID from $\{0,\dots,|G|-1\}$:
-  • Current timestep, $t\in T$
-  • ID the current node, $n_{t}$ from a list of all nodes in the environment graph
-  • A list with each index corresponding to a node ID. The element at a given index is the number of objects at the corresponding node at timestep $t$
-Action Space: The PPO agent then outputs the index of the node it wants to travel to for the next timestep $t+1$.
-We use a Maskable PPO to only allow it to choose neighboring nodes.
-The mask in this case is a list with each index corresponding to a node ID.
-The element at a given index is either $1$ if it is a neighbor of the current node and 00 otherwise.
-Reward Objective: Let $O_{t-1}$ be the set of distinct portable objects the agent as seen up to time $t-1$.
-The reward is the number of new objects it has found at the current node $n_{t}$ that are not in $O_{t-1}$.
-Formally, the reward at timestep $t$ is $|O(n_{t})\setminus O_{t-1}|$.
-Note. Our version of PPO uses a map of the environment and the number of objects at each node at each timestep.
-LLM-based Agents. For LLM-based navigation, we consider two variations of LGX [18], an LLM-based method for zero-shot embodied exploration.
-1.
-LLM-Vanilla: Given a target object, the LLM takes a list of objects around the agent and asks an LLM (GPT-4o) to predict an object to navigate toward.
-The agent then navigates towards the predicted object and continues doing so till the provided target is found.
-The LLM’s prompt contains a set of portable targets remaining to be found in the environment.
-At each timestep, we use YOLO v8 [19] to obtain a list of objects in the scene, and then prompt GPT-4o with the object list asking for an object prediction from the list that might lead to an unseen portable object.
-We then use the MP3D topological graph to hop to the next node.
-2.
-LLM with Memory: The vanilla version of the LLM agent works in a zero-shot manner resembling a Markov Chain, meaning that it makes navigation decisions at each node solely based on the observations it has seen at the current timestep.
-Since objects keep transiting on our OTG, keeping track of previously seen objects should help the agent make more prudent navigational decisions.
-We accommodate this by incorporating memory by passing a set of historical objects seen, predictions made, and timesteps to the prompt across episodes.
-The system prompt contains this historical data, along with a set of portable objects remaining to be found in the environment.
-We maintain a horizon to avoid GPT token overflow, where we remove the earliest appended observation when an overflow of tokens occurs.
-6 Experiments and Results
-We treat navigating in dynamic scenes and visual target grounding as separate problems in our experiments.
-This decoupling allows us to individually analyze navigation and grounding performance.
-The results of our visual grounding experiments are explained in detail in the supplementary.
-6.1 Hyperparameters
-We determine 4 hyperparameters that influence performance as described in the definition section, i.e., $S=[r,d,T,\Lambda]$.
-Each agent is subjected to a total of $[Stationary(1)+Transit(3)]\times R\times E\times T$ experiments, where $R$ is a set of rooms that varies with each scan from which a random node $r\in R$ is picked as a starting point for each trial.
-The maximum number of trials is thus equal to the number of rooms.
-$R$ ranges from $[11,30]$ in our chosen subset of $10$ MP3D scans.
-$E$ is the number of episodes that we run for each starting node $r$, and $T$ is the number of timesteps per episode.
-We set $R=10$, $E=20$ and $T=30$ in our experiments.
-Although different objects can have different levels of transition entropy, we set all objects in an experimental run to have the same transit scenario $\Lambda$.
-Setting all objects to have the same transit allows us to better analyze the generalizability of a navigation scheme in different levels of entropy in the OTG.
-Ground Truth Paths ($\Pi^{*}$).
-Given a starting node $r\in R$, for each episode $e\in E$, there exists a set of ground truth paths $\Pi^{*}$ over $T$ that the agent could take for collecting the most number of portable targets.
-We compute these paths by simulating all possible trajectories from a given starting point $r$, and store $\Pi^{*}$ to calculate performance and efficiency later.
-Object Density $d$.
-To make the MP3D environment less sparse with objects, we define a local graph of all nodes within $3$ steps away from the agent’s initial node.
-We then populate the local graph by adding duplicate portable objects from $O$ to the local graph.
-We populate until the object density $d$ of this local graph is $1$.
-As in, on average, each node in the local graph has $1$ object.
-In Algorithm 1, $d$ and the local graph size $q$ are hyperparameters.
-6.2 Evaluation Metrics
-Success Rate (SR): Given an episode, we measure the object-finding performance of an agent by dividing the number of objects found in that episode by the maximum possible number of objects that could have been found with a ground-truth path $\pi^{*}\in\Pi^{*}$: $SR=|O_{\pi}|/|O_{\pi^{*}}|$.
-This metric is similar to the Progress metric in Wani et al.
-[50].
-In this work, the ground-truth trajectory may not necessarily find all the objects in the given timeframe.
-Trajectory Alignment (TA): To measure path efficiency, we define Trajectory Alignment as the average overlap between ground-truth trajectory $\pi^{*}$ and trajectory taken by the agent across episodes.
-We take the maximum across $\Pi^{*}$: $TA=\min_{\pi^{*}\in\Pi^{*}}\frac{1}{T}\sum_{t\in T}\mathbf{1}_{\pi(t)=\pi^{*}(t)}$, where $\mathbf{1}$ is an indicator function.
-Relative Change in Success (RCS): We use this value taken from [17] to measure the generalizability from static to dynamic environments.
-It is the percent relative change between static and dynamic navigation performance.
-We measure RCS for various $\Lambda$ cases and report scores on SR and SPL.
-Lower values indicate better generalizability, showing that the agent performance between static and dynamic environments is consistent.
-Figure 4: Navigation Generalizability on OTGs (RCS): Relative Change in Success (RCS) measures an agent’s adaptability, comparing dynamic to static TG performance (optimal RCS is $0\%$).
-Positive values indicate better dynamic performance (chance encounters), while negative values reflect poor adaptability.
-Experiential-learning methods (RL, LLM+Memory) approach optimal RCS, demonstrating better generalization, while the Oracle and Vanilla LLM perform poorly.
-Random $\Lambda$ Semi-Routine $\Lambda$ Fully-Routine $\Lambda$ Avg. OTG
-Policy Metric (%) Static OTG RCS (%) Static OTG RCS (%) Static OTG RCS (%) Static OTG RCS (%)
-Random Agent SR $66.0$ $74.4$ $+11.2$ $66.0$ $94.7$ $+30.3$ $66.0$ $91.6$ $+27.9$ $66.0$ $86.9$ $+24.1$
-TA $33.5$ $24.4$ $-37.3$ $33.5$ $36.6$ $+8.4$ $33.5$ $33.6$ +0.29 $33.5$ $31.2$ $-6.8$
-Oracle Agent SR $90.2$ $78.4$ $-13.0$ $90.2$ $75.1$ $-16.8$ $90.2$ $55.9$ $-38.0$ $90.2$ $69.8$ $-22.6$
-TA $45.6$ $54.4$ $+16.2$ $45.6$ $50.6$ $+9.8$ $45.6$ $32.7$ $-28.3$ $45.6$ $45.9$ +0.6
-RL Agent SR $59.1$ $55.2$ $-6.6$ $59.1$ $57.8$2.3 $59.1$ $55.2$ $-6.7$ $59.1$ $56.0$ $-5.2$
-TA $41.6$ $37.3$ $-10.3$ $41.6$ $35.8$ $-14.0$ $41.6$ $32.1$ $-22.9$ $41.6$ $35.1$ $-15.7$
-Vanilla LLM Agent SR $28.5$ $43.3$ $+34.1$ $28.5$ $47.2$ $+39.6$ $28.5$ $34.2$ $+16.7$ $28.5$ $41.6$ $+31.4$
-TA $24.3$ $30.7$ $+20.9$ $24.3$ $31.3$ $+22.3$ $24.3$ $21.7$ $-10.8$ $24.3$ $27.9$ $+12.9$
-LLM + Mem. Agent SR $50.6$ $49.4$2.4 $50.6$ $53.2$ $+4.8$ $50.6$ $47.9$5.4 $50.6$ $50.2$0.9
-TA $36.9$ $33.8$8.5 $36.9$ $34.6$6.3 $36.9$ $32.4$ $-12.2$ $36.9$ $33.6$ $-9.0$
-Table 3: Navigation on OTGs: We evaluate $5$ popular embodied navigation approaches on various OTGs, and also on a static graph.
-These values are used to compute RCS using the equation provided in [17].
-An RCS value closer to $0\%$ indicates good generalizability to new conditions.
-Highlighted in green are the best RCS values for SR, while ones in violet are the best values for TA.
-Observe that the LLM + Memory agent shows the best SR generalizability overall, followed very closely by the RL agent.
-Both these agents employ experiential learning to learn about object transit patterns and improve their performance.
-6.3 Navigation Inference
-Table 3 presents the results of our agents on the navigation task on OTGs.
-We run experiments on $10$ Matterport3D scans, with $10$ randomly selected starting nodes $r\in R$, $20$ episodes, and each episode having $20$ timesteps.
-There are $4$ key inferences from these observations to understand the generalizability of the various navigation approaches:
-Experiential Navigation Approaches Generalize Well: Figure 4 presents a comparison of all approaches on RCS.
-LLM + Memory and PPO both have a relatively low RCS score, meaning that navigation agents perform comparably in stationary graphs and OTGs.
-Both methods rely on past observations to improve their decision-making, showing that data-driven approaches can be generalizable well.
-In contrast, the greedy heuristic Oracle agent and the zero-shot Vanilla LLM lack experience or memory, with the former relying on privileged knowledge and latter utilizing learned commonsense cues for decision-making.
-This result highlights learned decision-making improves generalizability, and we can infer this to be a key component for the real-world deployment of embodied agents.
-Zero-shot Methods Camp to Find Objects: In the Vanilla LLM approach which is zero-shot without memory, we observe that the agent finds more objects in the OTG than in the static environment, as clearly indicated by the higher OTG SR scores across all $\Lambda$s.
-In Table 3, we see that in the Fully-Routine case for Vanilla-LLM, the RCS for SR is $+16.7\%$ but the RCS for TA is $-10.8\%$.
-These results suggest that while Vanilla-LLM is finding more objects in OTGs than in static environments, its trajectories show lesser overlap with the ground-truth paths ($\Pi^{*}$).
-Upon inspection, we observe that this agent has learned to camp at privileged nodes that are in the paths of many object routes.
-One possible explanation is that the LLM is relying solely on common-sense knowledge and to go to areas where most objects can be found, rather than trying to learning object routes.
-We can consider camping as an instance of reward hacking, where we would not want an agent to camp at one location in anticipation of encountering a target.
-This challenge opens up exciting research directions in reward design for navigation on OTGs.
-Greedy Heuristic Struggles With Shifting Objects: The navigation approach of the agent heavily relies on the static placement of objects to achieve high performance.
-This case is best shown in the Oracle agent.
-At each timestep, the oracle takes a step towards the closest target.
-However since targets are on the move, a greedy oracle tends to take a suboptimal path.
-The agent is heading in one particular direction towards a certain target object which was nearest, but then suddenly switches to another direction due to the nearest unseen object changing.
-Entropy Influences Performance: In all policies barring the Random agent baseline, we observe that the performance on a Fully Routine $\Lambda$ exhibiting low entropy is usually worse than the Semi-Routine or Random $\Lambda$ cases.
-This is intuitive, and can be attributed to the agent encountering more targets in environments with higher object transit entropy.
-Despite this, we observe a strong performance of the RL agent in being entropy-agnostic, with very similar values across all transit cases on SR and SPL.
-7 Conclusion, Limitations, and Further Work
-We present Object Transition Graphs or OTGs, a generalization of the topological graphs to dynamic environments with shifting objects.
-Unlike static topological graphs where an agent is expected to navigate to stationary target objects, OTGs are realistic and challenging, with multiple portable objects transiting on a topological graph.
-Our dynamic task fundamentally challenges the static-scene assumption common in embodied navigation literature.
-We first introduce a novel approach to transform topological graphs to OTGs by developing structured object transition scenarios inspired by human habits.
-Our approach can be used to transform any topological graph into an OTG, and we transform graphs extracted from MP3D for this purpose.
-On these OTGs, we measure the generalizability of various navigation approaches performing a multi-object finding task.
-In comparing performance on static TGs and OTGs, we infer the generalizability of these approaches to dynamic settings with non-stationary targets.
-We find that the type of navigation employed greatly influences agent performance, with experiential approaches that have the means to learn from memory generalizing the best.
-In one case, we notice an instance of reward hacking via camping behavior, motivating future research in reward design for OTGs.
-Code and data, including the dynamized MP3D dataset, will be made publicly available to foster research in this exciting new direction.
-A limitation of our current setup is that we independently treat visually grounding small target objects as a separate problem.
-Coupling navigation on OTGs with visual grounding is a future direction.
-Another interesting direction in this regard is to utilize generative image modeling with OTG navigation on 2D simulators such as Minigrid [13], allowing us to scale up experimentation and training.
-Finally, our research paves the way for generalizability experiments in the context of dynamic scenes, and future work will look at studying the sim2real transfer of navigation policies trained on OTGs to a real world environment.
-Supplementary Material
-Supplementary Material
-8 Dataset Details
-While the agents introduced in Section 2 tackle the temporal dynamism, i.e., identifying where and when the target object is likely to lie at a particular location, visual grounding is a separate task tackling spatial dynamism.
-An ideal agent performs both these tasks simultaneously, verifying at each timestep if a portable target exists by running a grounding model.
-As navigation decisions are not directly impacted by the performance of a grounding model, we study temporal and spatial dynamism independently.
-The dataset contains temporal and spatial modifications on the Matterport3D scans. Further, we provide code and useful tools for implementing these modifications.
+8 DOMifying MP3D
+We convert MP3D graphs into DOMs with the $21$ different objects presented in Table 3.
+On each of the $10$ scans chosen, we introduce routine, semi-routine and random object movement, and provide these as dictionaries with varying seeds.
+We will release these modified environments along with supporting code.
+Table 4: Summary statistics of scans used for Table 2, including each scan’s ID, number of nodes, edges, and rooms.
+If a scan has multiple rooms of the same type (e.g.
+two bathrooms), each instance is counted separately from the total.
 Scan ID # of Nodes # of Edges # of Rooms
@@
 Z6MFQCViBuw 58 91 18
-Table 4: Summary statistics of scans used for Table 3, including each scan’s ID, number of nodes, edges, and rooms.
-If a scan has multiple rooms of the same type (e.g.
-two bathrooms), each instance is counted separately from the total.
-For the spatial placement, we provide a dataset of $10,500$ modified images taken across $10$ randomly chosen Matterport3D scans, with $21$ portable objects randomly oriented and positioned $5$ times within each image.
-Table 4 summarizes the different Matterport3D scans chosen.
-To emulate realism such that the portable objects are not floating in the sky or on the ceiling, they are placed on a large segmented area below the center of the image.
-We also provide the bounding boxes pertaining to the portable object in each image.
-8.1 Temporal: Implementing Object Placement Strategies
-In this section, we give more details on how we implemented the dynamic Matterport3D environment.
+8.1 Implementing Object Placement
+In this section, we give more details on how we convert Matterport3D environments into DOMs.
 Matterport3D Modifications:
@@
 Strategy Overview:
-We compute trajectories of the portable objects in an offline manner.
+We compute trajectories of the portable objects offline.
 For a fixed random seed $s$ and for each portable object $o_{p}\in O_{p}$, we create a sequence of nodes from the graph.
@@
 The resulting structure is a nested dictionary representing the changing graph.
-The keys of the outer dictionary are the timesteps [$1,...,T$].
+The keys of the outer dictionary are the timesteps [1,,T1,...,T].
 The inner dictionary for each timestep $t$ has the node string ids as the keys and the values are the list of portable objects at that node at $T$.
 At each timestep $t\in[1,...,T]$ when the agent reaches a node, we simply use $t$ and the node id to retrieve the list of portable objects the agent is currently observing.
-8.2 Spatial: Implementing Visual Grounding
-In this section, we talk about our strategy for spatial placement of multiple portable target objects on MP3D scenes.
-After placing portable objects at various locations, we perform visual grounding using three Open-World Object Detection models for inference.
-We also show further ablations using VLMs.
-Generating Spatial Data:
-We randomly choose $100$ skybox images taken across the $10$ MP3D scans mentioned in the previous subsection.
-While selecting these images are not of the ceiling or the floor, and have objects pertaining everyday scenes.
-We then pick stock images of all the $21$ portable objects present in Table 1 of the main manuscript.
-These images are shown in Figure 10.
-(a) Portable Target: Bowl
-(b) Portable Target: Playing Cards
-Figure 5: OWL-ViT Detection Failures: Note the poor detection accuracy of OWL-ViT.
-On both images, it completely misses the object when prompted with a text containing the target label.
-We attribute these failures to unnatural images being generated by pasting target objects onto the scene.
-(a) Portable Target: Playing Cards
-(b) Portable Target: Dumbbells
-Figure 6: GLIP Detection Accuracy: Note the superior detection accuracy, despite the low coverage on GLIP.
-On the image on the laft and right, playing cards and dumbbells are detected correctly with a high accuracy.
-(a) Portable Target: Screwdriver
-(b) Portable Target: Toothbrush
-Figure 7: YOLO Detection Failures: Note the poor detection accuracy of YOLO.
-On the image on the left, the screwdriver is grounded correctly, but the object name predicted is “Cow”.
-On the image on the right, YOLO completely misses the target object toothbrush, and instead grounds and labels something else.
-For each of the skybox images, we first run FastSAM [56] and segment out the largest regions below the center of the image.
-We then randomly orient the portable object and place it at a random position on one of these segments.
-Choosing segments below the center of the image is necessary to ensure that the portable objects being placed follow commonsense, and do not hang in the ceiling.
-This is done $5$ times for each of the $21$ objects on $100$ images, to get $10500$ images in total.
-While placing the images, we also store the boundaries of the overlay as the ground truth bounding box for computing Intersection over Union (IOU) scores.
-Examples of this generated data is shown in Figure 8.
-Note that the same image can have multiple target segmentations (bed and cushion for example), with different objects placed on them at different orientations.
-(a) Dice on Bed
-(b) Mug on Cushion
-(c) Flashlight on Floor
-(d) Charger on Table
-Figure 8: Various Spatial Placement in our Dataset: Figures (a) and (b) look at different arrangements of objects in the same room.
-Figure (c) shows a flashlight on the floor, and figure (d) shows a charger on the table.
-We cover various orientations and rotations of portable objects being placed on various objects in different MP3D scans.
-(a) bowl
-(b) dice
-(c) dumbbell
-(d) flashlight
-(e) glasses
-(f) hat
-(g) salt and pepper shakers
-(h) screwdriver
-(i) toothbrush
-(j) remote
-(k) playing cards
-(l) phone
-(m) phone charger
-(n) notebook
-(o) mug
-(a) laptop
-(b) jumprope
-(c) headphones
-(d) wallet
-(e) water bottle
-(f) wristwatch
-Figure 10: Portable Objects: A list of all the 21 portable objects we consider for our task.
-Open-World Visual Object Grounding
-We perform Visual Grounding on the spatial placement dataset using various Open-Vocabulary Object Detection Models.
-For each model, we assess the Detection Coverage, which is the percentage of images where the portable object was found, the Detection Accuracy, the accuracy of the predicted detections, and finally Mean IOU, which gives us the overlap between the ground truth bounding box of the placed target and the predicted box.
-The results for this are shown in Table 5.
-Approach Detection Coverage Detection Accuracy MIOU
-OWL-ViT[30] 100% $3.04\%$ $0.352$
-GLIP [24] $7.11\%$ 100% 0.489
-YOLO-World [12] $32.94$% $0.2$% $0.130$
-Table 5: Comparison of Open-World Object Detaction Approaches
-We make some interesting observations.
-OWL-ViT [30] outperforms GLIP [24] and YOLO-World [12] when it comes to Detection Coverage, which means that it consistently detects an object, irrespective of whether it is right or wrong on the image.
-This is evident since it takes a target object label is given as a prompt to OWL-ViT, so it detects something in the image, even if the confidence is low.
-Figure 5 shows some examples of this.
-GLIP, while having lower detection coverage, has a perfect detection accuracy, meaning that when GLIP does detect something, it usually is correct.
-The high MIOU score associated with GLIP also corroborates with this fact.
-Figure 6 shows some example results of this.
-We infer the YOLO-World model with custom vocabulary consisting of all the portable object names.
-Despite this, YOLO still performs the worst among the three, with poor detection accuracy and coverage.
-Figure 7 showcases one of these results.
-Notice that in the first image screwdriver is grounded at the right location, but the label is “Cow”.
-In the second image, both the predicted target (fire hydrant) and the grounding is wrong.
-We note that the poor detection as measured by MIOU could be a result of unnatural scenes that are different from those which these models might have been trained on.
-Beyond this, grounding small, portable targets that could be present on various surfaces in a household environment still remains a challenge.
-We will be releasing this dataset to foster research in this area.
-(a) Phone in bathroom
-(b) Headphones on bed
-Figure 11: We evaluate the spatial reasoning capabilities of VLMs in grounding small portable targets in houses.
-The question we ask is “What is the closest number to the $<$INSERT OBJECT$>$ in this image?” and it answers with a number.
-8.3 Spatial Reasoning with Vision Language Models (VLMs)
-We perform visual grounding using 4 popular Vision Language Models (VLMs): InstructBLIP [28], LLaVA 1.5 [27] and GPT-4o [1].
-For InstructBLIP, we experiment with two different language backends: Flan-T5, which has been finetuned on various language tasks including question answering, and Vicuna-13B, which has been trained specifically on conversational data.
-Figure 11 illustrates the images we pass to these models to validate their spatial reasoning capability.
-Given an image containing a small portable target for grounding, we first paste a number grid from $[1-9]$ onto the image.
-Prior research by Sathyamoorthy et.
-al [42] and Nasiriany et.
-al [33] has utilized this technique to guide robots in indoor as well as outdoor scenarios.
-In our case, we evaluate a subset of $1000$ images in our dataset containing various objects by asking each image the query - “ Where is the $<$objects$>$ in this image?
-Reply with the corresponding grid number from [1-9].”, and then evaluate the response we get with the ground truth.
-Since we create the dataset by pasting target objects onto them, this is readily available to us.
-Table 6 presents the accuracies for this task the various VLMs.
-Model Accuracy (%)
-InstructBLIP + Vicuna-7B 15.38
-InstructBLIP + Flan-T5-XL 4.32
-LLaVA v1.5 + Vicuna 13B 58.62
-GPT 4o-mini 75.89
-GPT 4o 98.32
-Table 6: VQA Model Accuracies: Note the superior performance of GPT 4o on our number grid based spatial reasoning task.
-The remaining VLMs show inferior performance, with a Vicuna backend performing better than the Flan one; this can be attributed to Vicuna being predominantly trained on conversational data that might contain such queries.
-9 Experimental Details
-9.1 LGX Inference
+8.2 TAS-PPO
+We modify a DD-PPO agent with temporal information about target objects in its observation space. As mentioned in section 3.3, we gather object paths $\mathcal{P}$ for augmenting this space.
+In our task, since we treat each object equally, i.e.
+the objective is simply to maximize finding targets, we include $\mathcal{P}$ as a single list $l=[v_{1},v_{2},\dots,v_{n}]$ of size $n=|G|$.
+Each index of $l$ corresponds to a different node in $G$.
+For each index $i$ of $l$, $v_{i}$ is the number of unique target objects at the $i$-th node in $G$.
+$l$ is updated at each timestep, and passed to the observation space along with the agent’s current position and timestep.
+The agent’s action space is masked, with the entire graph being the action space, and the mask allowing only for selecting neighboring nodes.
+Additionally, we swap out the distance reward $R_{dist}$ with the intersection reward $R_{I}$ described in the main text.
+$R_{I}$ here is triggered whenever the agent encounters a node where a target object has been in the last $5$ timesteps.
+We do not make any other modifications to the default setup.
+8.3 TAS-LLM
 We utilize the official implementation of LGX  [^1] https://github.com/vdorbala/LGX and incorporate it into our modified MP3D environment.
@@
 System Prompt - “I am a smart robot trying to find as many portable objects as I can at home.”
-User Prompt - “Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
-The $<$OBJECT_LIST$>$ here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX.
+User Prompt - “Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
+The <<OBJECT_LIST>> here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX.
 Additionally, we if a portable object is present at a certain node at a given timestep, we add it to this list.
-The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions.
-For the memory-enhanced LGX case, we modify the System Prompt, with memory. We additionally ask -
+The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions. The agent then moves to that node.
+For the TAS-LGX case, we modify the System Prompt, with memory as presented in Section 3.3. We additionally ask -
 System Prompt:
 “I have seen the following objects and taken the following actions so far -
-1. $<$OBJECT_LIST$>$: ACTION
-2. $<$OBJECT_LIST$>$: ACTION
+1. <<OBJECT_LIST>>: ACTION
+2. <<OBJECT_LIST>>: ACTION
 User Prompt:
-“Which object from $<$OBJECT_LIST$>$ should I go towards to find a new portable object? Reply in ONE word.”
-10 Code and Dataset
+“Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
+Note that we maintain a horizon to avoid token overflow, where the earliest appended observation is removed when the context length limit is reached.
+9 Video, Code and Dataset
 We provide an anonymous link to our Code: https://anonymous.4open.science/r/otg-1AC0. The dataset will be present as a downloadable link in the code repository.